### v1

# https://vl2g.github.io/projects/IndicPhotoOCR/

Abstract
Reading scene text, that is, text appearing in images, has numerous application areas, including assistive technology, search, and e-commerce. Although scene text recognition in English has advanced significantly and is often considered nearly a solved problem, Indian language scene text recognition remains an open challenge. This is due to script diversity, non-standard fonts, and varying writing styles, and, more importantly, the lack of high-quality datasets and open-source models.
To address these gaps, we introduce the Bharat Scene Text Dataset (BSTD) – a large-scale and comprehensive benchmark for studying Indian Language Scene Text Recognition. It comprises more than 100K words that span 11 Indian languages and English, sourced from over 6,500 scene images captured across various linguistic regions of India. The dataset is meticulously annotated and supports multiple scene text tasks, including: (i) Scene Text Detection, (ii) Script Identification, (iii) Cropped Word Recognition, and (iv) End-to-End Scene Text Recognition. We evaluated state-of-the-art models originally developed for English by adapting (fine-tuning) them for Indian languages. Our results highlight the challenges and opportunities in Indian language scene text recognition. We believe that this dataset represents a significant step toward advancing research in this domain. All our models and data are open source.
keywords:
Multilingual Scene Text Recognition, Bharat Scene Text Dataset, IndicPhotoOCR
Figure 1: Scene Text Understanding: Case of India. This figure showcases typical street scenes from northern, eastern, western, and southern parts of India, highlighting the country’s vast linguistic diversity. With 11 languages (five of them: Hindi, English, Bengali Gujarati and Tamil are shown here) including English commonly featured on signboards, scene text understanding in India presents unique challenges. While significant progress has been made in English Scene Text Recognition, open-source comprehensive effort for Indian language scene text understanding, including large public datasets and comprehensive models, are still limited. Our work seeks to address this gap and advance the field.
1 Introduction
Scene Text Recognition, especially for English, has advanced significantly in the last decade, thanks to a series of efforts on dataset creation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and model development [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In contrast, scene text recognition for Indian languages has been extremely underexplored and possibly more challenging (refer to Fig. 1). Note that Indian languages are used by more than 1.4 billion people, which is approximately 18% of the global population. Despite its usage and popularity, we observe that scene text recognition in Indian languages lacks a comprehensive benchmark and open-source models, except for some early efforts [26, 27, 28, 29, 30, 31, 32, 33] that laid important groundwork, but did not aim for a comprehensive coverage of tasks or languages. We aim to fill this gap through our work.
To accelerate advancements in Indian Language Scene Text Recognition, we present a comprehensive benchmark called the Bharat Scene Text Dataset (BSTD) [^1] India, traditionally known as Bharat, reflects the nation’s rich cultural and linguistic diversity, which inspired the naming of our dataset.. The BSTD comprises 6.5K scene images that feature more than 100K words in 11 Indian languages, along with English. These images, sourced from Wikimedia [^2] https://www.wikimedia.org/, span various signboards and accidental background text from different Indian cities and towns, ensuring that the dataset captures the linguistic diversity of India. We obtain a polygon-level bounding box for each visible text in the image and the text it contains using rigorous manual annotation. The dataset is publicly available for download here [^3] https://github.com/Bhashini-IITJ/BharatSceneTextDataset.
The BSTD can be used for the following scene text tasks: (i) scene text detection, (ii) script identification, (iii) cropped word recognition, and (iv) end-to-end scene text recognition. In this work, we benchmark the BSTD using our new pipeline, [Uncaptioned image], which integrates state-of-the-art architectures across its modules. We firmly believe that these benchmarks will help accelerate advancements in Indian language scene text recognition.
Dataset # Images # Words # Languages Tasks Supported
Detection Script Identification Recognition
ICDAR 2003 [1] 509 2,268 1
SVT [2] 350 725 1
IIIT-5K [3] - 5,000 1
ICDAR 2013 [5] 462 5,003 1
ICDAR 2015 [6] 1,500 6,545 1
CUTE80 [7] 80 288 1
Total-Text [8] 1,555 11,459 1
IIIT-TL-STR [30] 440 3,035 2
IIIT-ILST [27] - 3,168 3
MLT-17 [34] 18,000 96,000 9
MLT-19 [10] 20,000 191,000 10
IIIT-indicSTR12∗∗ [31] - 27,000 12
BSTD (This Work) 6,582 1,06,478 12
Table 1: Comparison of the Bharat Scene Text Dataset (BSTD) with existing scene text datasets. BSTD supports all three core tasks—text detection, script identification, and recognition (and, thereby end-to-end evaluation) across 12 Indian languages, including English. It is the only publicly available dataset for Indian languages that enables comprehensive end-to-end scene text understanding. : MLT-17 and MLT-19 are not focused on Indian languages, they only include Hindi and Bengali. **: IIIT-IndicSTR12 is primarily annotated for cropped word recognition and can not be studied to evaluate end-to-end scene text recognition as ours.
Figure 2: Language map of India (Source: Wikimedia). The BSTD covers scene text images from 11 prominent languages namely Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Punjabi, Tamil and, Telugu. Additionally, it contains English, as English is often included as one of the languages as signboard in India.
We list our contributions here:
  1. 1.
    We introduce the Bharat Scene Text Dataset (BSTD) – a large-scale benchmark for Indian language scene text recognition, consisting of 6,582 scene images featuring 1,06,478 words in 11 Indian languages and English. Each image is manually annotated with polygon-level bounding boxes and corresponding text and script, ensuring high-quality data for research and applications. (Section 2)
  2. 2.
    We evaluated BSTD using state-of-the-art techniques, including off-the-shelf methods and fine-tuned models on synthetic and/or BSTD train data (as applicable). Detailed benchmark results are provided for all key tasks, i.e., text detection, script identification, cropped word recognition, and end-to-end scene text recognition. These evaluations establish a strong baseline for future research and advancements in multilingual scene text recognition, particularly for Indian languages. (Section 4)
  3. 3.
    We introduce an open source toolkit, namely [Uncaptioned image], designed to detect, identify, and recognize text in English and 11 Indian languages. The toolkit is easy to install and replicate, making it accessible to researchers and developers. (Section 5)
2 Bharat Scene Text Dataset
India, often referred to as Bharat (the Indian name for the country), represents a rich and diverse cultural and linguistic landscape, as illustrated in Fig. 2. In the context of scene text understanding, the vast multilingual environment poses unique challenges, as the scripts, writing styles, and text formats vary significantly between different regions. In addition, the identification of scripts becomes extremely important. Although there has been some effort in collecting and benchmarking Indian language scene text understanding tasks, it falls short compared to efforts for scene text understanding in English [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Please refer to Table 1 for a detailed comparison of the Bharat Scene Text Dataset compared to other related scene text datasets.
To address this gap, we present the Bharat Scene Text Dataset, a novel and comprehensive dataset specifically curated to represent the linguistic diversity of India. By using the term “Bharat”, we not only acknowledge the nation’s deep-rooted cultural identity but also highlight our focus on supporting scene text understanding across 11 major Indian languages. This dataset contains 6,582 scene images and 1,06,478 word images in total. This dataset and associated annotations can be downloaded from our project website [^4] https://vl2g.github.io/projects/IndicPhotoOCR/. In this section, we give details of dataset curation and annotation, supported tasks, train-test splits followed by dataset analysis.
2.1 Dataset Curation and Annotation
We illustrate the pipeline for constructing the Bharat Scene Text Dataset from Wikimedia Commons in Fig. 3. The dataset is curated and annotated in the following three stages:
Figure 3: Pipeline for constructing a multilingual scene-text dataset from Wikimedia Commons images. The process begins with query construction using combinations of Indian city names and place-related keywords to retrieve relevant images via the Wikimedia Commons API. A two-stage filtering process using DBnet [19] (pre-trained on Indic Synthetic Data) removes non-textual images and manually filters relevant ones. Scene-text detection annotations are generated using DBnet and refined through manual corrections. Finally, cropped word images are recognized using the PARSeq [22] model (trained on 11 Indic languages using synthetic data), with recognition errors corrected and script language tags added to produce the final annotated dataset.
2.1.1 Stage-1: Image collection
We compile a comprehensive list of cities and towns in various regions of India where 11 selected languages are spoken predominantly. Furthermore, we curate a catalog of public places, such as ‘bus stations’, ‘banks’, ‘schools’, ‘railway stations’, ‘hospitals’, ‘ATMs’, ‘metro stations’, ‘street signs’, ‘airport’, ‘shop entrance’, ‘billboards’, ’press conference’, etc. By combining the city and place names as search queries, e.g. ‘Delhi Railway Station’, we retrieved images along with their metadata from the Wikimedia Commons API [^5] https://commons.wikimedia.org/w/api.php. It should be noted that not all retrieved images contain scene text at this stage, which requires subsequent filtering. At this stage, we obtain 50K images.
Figure 4: An example image from BSTD along with the corresponding JSON annotation.
2.1.2 Stage-2: Image filtering
We filter the images harvested in the previous step using the pre-trained scene text detection method [19]. To this end, we run this detector and obtain text even with low confidence. We retain images containing at least one word, while others were discarded. In addition, we ask human annotators to review the filtered set to eliminate false positives. The final data set comprises 6,582 images that contain scene text in 11 Indian languages and English.
2.1.3 Stage-3: Semi-automatic annotation
We annotate the final filtered images for scene text detection in two steps to reduce human annotation efforts. We obtain bounding box annotation using a fine-tuned DBnet model on synthetic Indian language data [^6] Refer to Section-3.3.1 for details of synthetic data generation. Since these automatically obtained bounding boxes are not always perfect, they are refined by our hired annotators using an in-house tool to edit, draw, or delete polygons around the scene text words.
Once the bounding boxes were obtained, they were assigned to the annotators and assigned to label each box with the script it contains. These annotators were well-versed in script identification and were also informed about the geographical origin of each image. With this context, they were able to accurately label each bounding box with the corresponding script. We then use PARSeq [22], fine-tuned on synthetic data for Indian languages, to read the text within each bounding box. The bounding boxes, along with the recognized text, are then provided to language expert annotators. These annotators review the text and correct any recognition errors or retain the text as is if it is accurate.
2.2 Annotator Details
We employed two distinct groups of annotators for this work. Initially, three human annotators for image filtering (Section 2.1.2) and bounding box annotation (Section 2.1.3). For the next two steps, text correction and script assignment (refer 2.1.3), we hired eleven language-specific experts, one for each target language. The annotators were proficient in reading, understanding, and writing their native language, as well as English. They were shown cropped words and asked to correct the pseudo-recognition annotations and assign the script. All final annotations underwent an additional quality check to eliminate potential errors and noise. The annotators were compensated at rates aligned with standard wages for a tier-2 city in India. The entire process–pseudo-annotation, manual correction, and quality verification took 10 months. Fig 4 shows the fine grained nature of the annotations done in this work.
2.3 Supported Tasks and Evaluation Protocols
The Bharat Scene Text Dataset (BSTD) enables us to study the following four tasks:
  1. 1.
    Scene Text Detection: where given an image, the task is to detect all the text instances in the image along with their bounding boxes. We provide polygon-level tight bounding boxes for each word image in the dataset. The popular metrics such as Precision, Recall and F1-score using TedEval evaluation [35] can be used to assess the performance of the method.
  2. 2.
    Script Identification: This task involves identifying the language of the script present in a given cropped word image. The performance of the model is evaluated using accuracy. Additionally, we include a confusion matrix to analyze where the model’s predictions fail.
  3. 3.
    Cropped Word Recognition: where the task is to recognize and extract the text given a cropped word image provided the script/language of text is known. For cropped word recognition, the word recognition rate (WRR) is used as a performance metric.
  4. 4.
    End-to-End Scene Text Recognition: This task is the closest to the natural setting. Here, given a scene image, the task is to recognize all the instances of text. Solving end-to-end Scene Text Recognition for Indian languages, often requires a pipeline of scene text detection, script identification followed by cropped word recognition. We use mean word recognition rate (WRR) and character recognition rate (CRR) over all the test images in the dataset. The WRR and CRR for an image are defined as follows:
    $\text{WRR}=\left(1-\frac{S_{w}+D_{w}+I_{w}}{N_{w}^{\text{total}}}\right)\times 100$ (1)
    $\text{CRR}=\left(1-\frac{S_{c}+D_{c}+I_{c}}{N_{c}^{\text{total}}}\right)\times 100$ (2)
    where $S$, $D$, and $I$ with subscript $w$ and $c$ denote the number of substitutions, deletions, and insertions, respectively, at the word or character level. Further, $N_{w}^{\text{total}}$ and $N_{c}^{\text{total}}$ denote the total number of words and characters present in the ground-truth transcription in the image, respectively. It should be noted that due to false alarms (i.e., excessive insertions of words or characters in the recognized text) or in some rare cases of any missing annotation, the computed WRR or CRR values can become negative for few images. To keep the interpretation simple, we truncate such values to 0. Furthermore, standard WRR and CRR computations assume a fixed reading order; however, since reading order in scene text images can be subjective, we relax this constraint in our evaluation. Further, to make interpretation of results more clean, we also use standard precision, recall, and F1-score to evaluate end-to-end performance. Note that precision is defined as the number of correctly recognized words over the total number of recognized words, while recall is defined as the number of correctly recognized words over the total number of words in the ground truth.
Figure 5: Box plot highlighting the word length range across all the 12 languages.
We randomly divided the 6,582 images into an 80-20% train-test split. As a result, the training set consists of 5,263 scene images and 77,726 cropped words, each with corresponding language and text annotations. The test set contains 1,319 scene images and 28,752 cropped words with their respective language and text annotations. The image-wise and cropped world-level train/test split is summarized in Table 2 and Table 3, respectively. For the script identification task, we randomly selected an equal number of word images from each of the 12 languages. The split comprises 1,800 training images and 478 testing images per language.
Set Images Bounding Box Annotations
Total 6,582 1,26,292
Train 5,263 94,128
Test 1,319 32,164
Table 2: Summary of image and detection annotation counts in the BSTD dataset, detailing the total, training, and test set distributions
Language #Word images Train Test
Assamese 4,132 2,627 1,505
Bengali 6,304 4,936 1,368
English 41,696 29,123 12,573
Gujarati 2,899 1,884 1,015
Hindi 19,773 14,927 4,846
Kannada 2,928 2,208 720
Malayalam 2,940 2,393 547
Marathi 5,113 3,917 1,196
Odia 4,192 3,148 1,044
Punjabi 11,199 8,319 2,880
Tamil 2,542 2,029 513
Telugu 2,760 2,215 545
Others 19,814 - -
Total 1,26,292 77,726 28,752
Table 3: Cropped word image data split for the BSTD dataset, showing the number of annotated instances, train images, and test images across multiple languages. Here, Others correspond to those words which are not from the listed languages here (e.g., Meitei and Urdu). We plan to include their annotation in subsequent versions of our dataset.
Figure 6: Distribution of the bounding box size in our dataset, in terms of pixels.
2.4 Dataset Analysis
In this section, we present a comprehensive analysis of the Bharat Scene Text Dataset (BSTD) to evaluate its quality and the challenges it poses. Toward this end, we first report the language-wise word count in Table 3. We observe that it has a good coverage of all the major Indian languages with minimum 2,542 word images for Tamil. This makes the dataset linguistically rich. We further analyze the average word length for each language in our dataset in Fig. 5). The plot reveals that Malayalam, Tamil, and Telugu contain longer words on average, whereas English and Hindi have shorter word lengths. This observation, though potentially intuitive, holds practical implications for the development of specialized language models in the future.
Moving further, we studied the distribution of the bounding box area in Fig. 6. It indicates that most text regions (59.4%) are small (1001–10,000 pixels), while only 2.6% of text instances fall into the largest category (above 100,000 pixels). These findings highlight the diverse textual characteristics within BSTD and it contains a mix of small and large text regions within images making it challenging for detection. Furthermore, the high percentage of small text regions suggests that many images contain incidental text rather than just focused text, which adds to the dataset’s difficulty implying the real word scenario.
Figure 7: Word cloud for recognition annotations in our BSTD dataset: Highlighting the diversity of vocabulary across 12 languages. The tiles are arranged row-wise from left to right as follows: the first row contains Hindi, Assamese, and Gujarati; the second row shows English, Malayalam, and Kannada; the third row corresponds to Odia, Panjabi, and Tamil; and the fourth row includes Telugu, Marathi, and Bengali.
Figure 8: A selection of challenging cropped words from our dataset. Few key challenges include eroded text (c, d, k), handwritten text (f), occluded text (e, h, l), ambiguous text (a, b, i), text in different orientations (g, j).
Finally, to illustrate diversity in the dataset, we show word clouds for all 12 languages in our data set in Fig. 7. As can be seen, the most frequent words (represented by the largest font size) across all languages point to text commonly found on public signs, government buildings, and commercial establishments. In addition, the vocabulary is dominated by nouns that are staples of scene text recognition tasks. These are words that one would expect to see “in the wild”. Some of the cropped word samples from BSTD samples are shown in Fig. 8.
In short, the analysis of the Bharat Scene Text Dataset suggests that it has diverse linguistic and geographical coverage, with a clear focus on text from public, commercial and transport-related signage, and has a mix of focused and incidental text, making it a valuable resource for developing robust real-world text recognition models for Indian languages. Further, while widely spoken languages such as Hindi, English, Bengali, and Marathi are well represented in the dataset, low-resource languages including Assamese, Gujarati, Kannada, Malayalam, Odia, Punjabi, Tamil, and Telugu also have a substantial presence, with at least 2,542 word images for Tamil (see Table 3). This shows that the dataset is not only strong in high-resource languages but also sufficiently representative of low-resource languages, making it suitable for comprehensive multilingual text recognition research.
3 [Uncaptioned image]: Proposed Baseline
Scene text recognition in a multilingual environment typically involves three key modules: (i) Scene Text Detection, (ii) Script Identification, and (iii) Cropped Word Recognition. In the absence of a dedicated architecture for Indian language scene text recognition in the literature, we develop a strong baseline by adapting or fine-tuning (as needed) state-of-the-art architectures across various modules. We describe each of these modules below and collectively refer to this pipeline as [Uncaptioned image].
3.1 Scene Text Detection
Our systematic analysis of modern scene text detectors revealed that these models are largely language-agnostic. Hence, we directly incorporate them into our pipeline without language-specific modifications. In particular, we adopt TextBPN++ [24] for scene text detection, as it demonstrated superior performance compared to other detectors in our evaluations. Detailed results for this stage are presented in the experimental section 4.1.
3.2 Script Identification
For script classification, we used the google/vit-base-patch16-224-in21k Vision Transformer (ViT) model, which consists of 12 transformer layers and uses $16\times 16$ patch embeddings. Pre-trained on the large-scale ImageNet-21k dataset, this model provides rich visual representations that are well-suited for transfer learning for the task of script identification.
To adapt the model for our task, we appended a custom classification head that projects the learned visual features into output logits, corresponding to either 12 classes (for full multilingual classification) or 3 classes (a practical setting with Hindi, English, and one regional language at a time).
We fine-tuned the model, including both the transformer backbone and the classification head, using the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a batch size of 16 using BSTD-train set. This ViT-based approach consistently outperformed other configurations, demonstrating strong generalization in both script classification settings of the 3-way and 12-way scripts.
Language Train Size Test Size Vocabulary Size #Fonts
Assamese 8,092,205 899,134 636,312 44
Bengali 8,475,805 941,757 4,447,622 26
Gujarati 6,431,896 714,656 4,255,185 57
Hindi 8,357,906 811,976 4,195,486 101
Kannada 4,002,547 444,728 8,288,718 22
Malayalam 7,199,420 400,000 4,857,185 32
Marathi 3,168,386 352,043 3,739,734 101
Odia 4,041,852 449,095 1,144,724 35
Punjabi 5,479,762 608,863 2,083,909 40
Tamil 5,652,718 628,080 7,842,880 70
Telugu 3,485,570 387,286 5,705,371 42
Table 4: Statistics for Synthetic Dataset generated using [36] for 11 Indian languages.
3.3 Cropped Word Recognition
Once text instances are detected and their scripts identified, the next step is text recognition. Unlike detection, this task relies heavily on language-specific modeling and requires large-scale training data. Off-the-shelf models for English text recognition cannot be directly applied as they are not designed for this task. Instead, we fine-tune a successful scene text recognition architecture (namely PARSeq [22]) for the 11 Indian languages in our study. A key challenge in this setting is the scarcity of real-world scene text data for low-resource languages, in contrast to English, where recognizers are trained on millions of samples. To address this, we generated synthetic word images to augment the training data and bridge this gap as described in the following section.
3.3.1 Synthetic Data Generation
To train models for each of the 11 languages, we required a large-scale dataset of scene text images. However, such datasets for Indian languages were not readily available. This lack of data led us to adopt an existing method for generating synthetic scene text images. We used SynthText [36] to generate large-scale data, with statistics summarized in Table 4. SynthText [36] provided 8000 background images, which we leveraged as the background image for our synthetic data.
To generate scene text in different Indian languages, we needed appropriate vocabulary and fonts for each language. We sourced vocabulary from the AI4Bharat-IndicNLP Dataset [^7] https://github.com/AI4Bharat/indicnlp_corpus introduced in the IndicNLPSuite [29]. SynthText also accounts for local 3D geometry, enabling varied orientation of text placement within the scene. Although the generated images lacked complete realism, the scale of data generation offered significant value for our training needs.
This synthetically generated data is then used to train the PARSeq model for each language. PARSeq requires accurate character sets, which vary by language (e.g. Assamese: 91, Bengali: 91, Gujarati: 89; Hindi: 125, Kannada: 85, Mayalayam: 116, Marathi: 125, Odia: 90, Punjabi:75, Tamil: 64, Telugu: 98). During this stage, we used an input image size of $32\times 128$, a maximum label length of 25, and random augmentation. Following synthetic training, the best model is then fine-tuned on real data (BSTD) using decaying learning rate applied layer-wise.
4 Experimental Analysis
In this section, we provide a detailed evaluation of the proposed baseline, namely [Uncaptioned image], and its individual modules on the newly introduced Bharat Scene Text Dataset. We compare it with several open-source OCR systems, including Tesseract [37], PaddleOCR [38], EasyOCR [39], and SuryaOCR [40]. For end-to-end evaluation, we also compare it with commercial closed-source systems such as Google OCR [41] and GPT-4 [42]. In addition, we include appropriate competitive approaches to evaluate individual modules. Our results reveal both the strengths and limitations of existing systems in handling diverse multilingual Indian scene text. They also demonstrate the effectiveness of [Uncaptioned image] in a challenging setting. Please note that all the results are reported on the Bharat Scene Text Dataset (test set).
4.1 Results on Scene-text detection
Our proposed baseline, [Uncaptioned image], employs TextBPN++ [24] for scene text detection. We compare its performance with several widely used scene text detection methods, including EAST [14], CRAFT [16], DBNet [19], and Hi-SAM [43]. The comparison results are summarized in Table 5. We observe that TextBPN++ significantly outperforms the other detectors, achieving higher F1-score and recall. Although Hi-SAM is also equally precise, it misses out many in detecting many Indian language words. Please note that we use TedEval [35] evaluation for reporting these numbers.
Method P R F1
EAST [14] 0.55 0.10 0.17
CRAFT [16] 0.51 0.22 0.19
DBNet [19] 0.71 0.50 0.59
Hi-SAM [43] 0.75 0.19 0.46
TextBPN++ [24] 0.75 0.78 0.77
Table 5: Scene Text Detection results on BSTD-Test Set using techniques prevalent in English scene text datasets.
4.2 Results on Script Identification
For script identification, we evaluate performance under two settings: (i) 3-way classification, where the model distinguishes among a regional language, Hindi and English. This setup reflects a common real-world scenario in India, where signboards often include all three languages. (ii) 12-way classification, where the model selects from the English and 11 Indian languages.
We use the following baselines: (i) AlexNet [44]: a popular CNN-based Imagenet-pretrained image classification model finetuned on trainset of our script identification data, (ii) CRNN [45]: a CNN-based scene text recognition model with the ability to encode sequence information, adapted for script identification, followed by state-of-the-art transformer architectures for image classification; (iii) CLIP [46] and (iv) ViT [47], respectively. All the results reported in Table 6 correspond to the models trained on real images.
Trained on BSTD
Language AlexNet [44] CRNN [12] CLIP [46] Proposed
Assamese 53.1 72 75.6 92.2
Bengali 54 75.7 77.5 93.3
Kannada 54.7 73.8 82.7 94.5
Malayalam 54.3 71.6 84.1 90.8
Marathi 45.9 62.3 70.1 79.6
Odia 50.1 87.2 86.7 93.5
Tamil 51.3 75.3 79.6 93.6
Telugu 54.9 82.9 86.9 95.0
Gujarati 50.7 70.2 79.1 83.1
Punjabi 51.3 81.1 79.7 92.3
Table 6: Script Identification Accuracy in 3-way classification setting on Hindi vs English vs third regional Indian language shown in the language column.
Figure 9: Normalized confusion matrix illustrating the performance of script identification model across 12 Indian. (Best viewed in color).
For the 3-way classification, the results are reported in Table 6. We observe that VIT-based proposed script identification clearly outperforms other baselines. For 12-way script classification, the CLIP-based baseline achieves an accuracy of 67.7%, while our proposed ViT-based approach achieves 80.5%. The corresponding normalized confusion matrix is presented in Fig. 9. While the overall results are satisfactory, the model frequently confuses Assamese with Bengali and Hindi with Marathi. This is primarily due to their sharing Unicode patterns. Also, since they share majority of the Unicodes, these two confusion does not affect significantly our end-to-end performance. We have also noticed that other languages frequently confuse with English, likely due to some data bias or weak context for small words (such as ‘&’, ’ON’, etc.). Our end-to-end evaluation suggests that further improving script identification can significantly boost overall scene text recognition performance.
4.3 Results on Cropped Word Recognition
For the cropped-word recognition baselines, there are no models specifically designed for Indian languages. As a result, we must rely on English scene-text recognition approaches such as [25, 48, 22, 45]. However, adopting all these methods still demands substantial computational resources, especially when running recognition models across 12 languages. Therefore, we only choose PARSeq [22], one of the recent and effective recognition models, as our baseline. In addition, we compare the PARSeq-based baseline with the already available, CRNN [12] by AI4Bharat [49] and off-the-shelf open-source models, namely Tesseract [37], PaddleOCR [38], and EasyOCR [39].
ParSeq is trained on a combination of synthetic data and the BSTD dataset. Our baseline for cropped word recognition is based on the existing text recognition framework PARSeq [22]. We report results under two settings: (i) training PARSeq from scratch using synthetic data, and (ii) fine-tuning it on the BSTD training set. The results are presented in Table 7. Fine-tuning on real data leads to substantial improvements in recognition accuracy. This highlights the critical role of real-world data in scene text recognition. Also, as can be seen, Word Recognition Rates for cropped English scene text are relatively high (around 92%). However, performance drops noticeably for other languages: Marathi (86%), Bengali (82%), Tamil (80%), Aasamease (79%) Punjabi (75%), Odia (72%), and Hindi (71%) form the next tier of top-performing languages. The gap widens further for low-resource languages such as Telugu and Malayalam, whose recognition accuracies remain in the 50% range. These results clearly indicate substantial room for improving recognition performance across Indian languages.
In the next section we will be evaluating recognition performance in even more challenging end-to-end setting.
Model Train Data Assamese Bengali English Gujarati Hindi Kannada Malayalam Marathi Odia Punjabi Tamil Telugu Average
PARSeq [22] Synth 0.47 0.44 0.92 0.32 0.43 0.47 0.52 0.44 0.42 0.4 0.37 0.44 0.47
BSTD Train 0.79 0.82 0.92 0.61 0.71 0.6 0.58 0.86 0.72 0.75 0.8 0.56 0.73
CRNN [12] AI4Bharat NA NA 0.27 NA 0.13 NA 0.06 NA NA 0.27 0.16 0.05 0.16
Off-the-shelf open source models
Tesseract [37] - 0.15 0.18 0.33 0.12 0.2 0.12 0.03 0.21 0.11 0.19 0.13 0.09 0.15
PaddleOCR [38] - NA NA 0.61 NA 0.31 0.03 NA 0.36 NA NA 0.28 0.15 0.29
EasyOCR [39] - 0.12 0.24 0.29 NA 0.18 0.06 NA 0.25 NA NA NA 0.11 0.18
Table 7: Baseline Results on Cropped Word Scene Text Recognition for 12 languages on BSTD. Please note that PARSeq in used in our proposed baseline – [Uncaptioned image].
WRR on BSTD-Test
Method /b Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
ChatGPT [42] 0.08 0.36 0.09 0.2 0.03 0.05 0.04 0.12 0.21 0.1 0.13
Google OCR [41] 0.43 0.37 0.32 0.55 0.24 0.37 0.43 0.46 0.52 0.42 0.41
Tesseract [37] b 0.04 0.02 0.02 0.03 0 0 0.02 0.02 0 0.02 0.02
SuryaOCR [40] b 0.18 0.15 0.09 0.19 0.07 0.12 0.15 0.21 0.24 0.05 0.14
[Uncaptioned image] b 0.45 0.37 0.3 0.39 0.29 0.31 0.38 0.41 0.42 0.31 0.36
[Uncaptioned image] (+oracle TD) b 0.75 0.74 0.54 0.63 0.49 0.54 0.68 0.7 0.69 0.49 0.62
[Uncaptioned image] (+oracle TD+SI) b 0.8 0.92 0.61 0.74 0.6 0.58 0.72 0.75 0.8 0.56 0.71
CRR on BSTD-Test
Method /b Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
ChatGPT [42] 0.13 0.46 0.15 0.28 0.1 0.14 0.11 0.2 0.31 0.21 0.21
Google OCR [41] 0.57 0.42 0.36 0.66 0.45 0.55 0.57 0.59 0.69 0.6 0.55
Tesseract [37] b 0.12 0.07 0.06 0.09 0.05 0.09 0.09 0.09 0.04 0.06 0.08
SuryaOCR [40] b 0.29 0.24 0.15 0.33 0.19 0.26 0.29 0.31 0.45 0.26 0.28
[Uncaptioned image] b 0.59 0.47 0.43 0.54 0.48 0.55 0.59 0.53 0.62 0.57 0.54
[Uncaptioned image] (+oracle TD) b 0.87 0.78 0.64 0.74 0.68 0.8 0.84 0.84 0.78 0.73 0.77
[Uncaptioned image] (+oracle TD+SI) b 0.93 0.97 0.75 0.89 0.86 0.86 0.89 0.9 0.94 0.83 0.88
Table 8: We conduct a comparative evaluation of commercial systems, including Google OCR and GPT-4, alongside our proposed baseline namely [Uncaptioned image] for end-to-end scene text recognition on the BSTD dataset. The evaluation measures script-wise performance using Word Recognition Rate (WRR) and Character Recognition Rate (CRR). Additionally, we present baseline results under two scenarios: (i) when text detection is provided as ground truth (denoted as oracle TD) and (ii) when both text detection and script identification are given as ground truth (denoted as TD+SI). “ /b” indicates whether the tool is open-source ( b) or closed-source (). Bold and underline number shows best performing results in non-oracle setting for closed and open-source models respectively. [Uncaptioned image] is the best open-source model for all the Indian languages. The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.
Method /b Metric Languages Avg.
Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu
ChatGPT [42] P 0.11 0.51 0.15 0.28 0.05 0.07 0.06 0.17 0.23 0.11 0.17
R 0.09 0.42 0.12 0.23 0.05 0.06 0.06 0.15 0.22 0.1 0.15
F 0.09 0.44 0.13 0.24 0.05 0.06 0.05 0.15 0.22 0.1 0.15
Google OCR [41] P 0.52 0.54 0.42 0.64 0.34 0.45 0.54 0.59 0.58 0.49 0.51
R 0.55 0.71 0.44 0.65 0.3 0.42 0.61 0.6 0.64 0.53 0.54
F 0.52 0.59 0.42 0.63 0.31 0.43 0.56 0.58 0.6 0.5 0.51
Tesseract [37] b P 0.08 0.05 0.05 0.06 0.01 0 0.04 0.08 0 0.03 0.04
R 0.08 0.11 0.03 0.05 0.02 0.01 0.04 0.06 0.01 0.02 0.04
F 0.07 0.06 0.03 0.05 0.01 0.01 0.03 0.05 0 0.02 0.03
SuryaOCR [40] b P 0.27 0.3 0.19 0.29 0.11 0.17 0.23 0.38 0.34 0.1 0.24
R 0.24 0.36 0.1 0.23 0.09 0.15 0.18 0.24 0.29 0.13 0.20
F 0.23 0.29 0.12 0.25 0.09 0.15 0.19 0.28 0.29 0.11 0.2
[Uncaptioned image] b P 0.59 0.55 0.43 0.53 0.41 0.39 0.47 0.56 0.55 0.38 0.49
R 0.53 0.46 0.4 0.43 0.33 0.38 0.47 0.49 0.52 0.37 0.44
F 0.55 0.48 0.4 0.46 0.36 0.38 0.47 0.51 0.52 0.37 0.45
[Uncaptioned image](+oracle TD) b P 0.72 0.83 0.65 0.66 0.56 0.57 0.65 0.69 0.78 0.54 0.66
R 0.69 0.75 0.62 0.61 0.5 0.57 0.63 0.68 0.73 0.53 0.63
F 0.7 0.78 0.62 0.62 0.52 0.57 0.63 0.68 0.75 0.53 0.64
[Uncaptioned image](+oracle TD+SI) b P 0.75 0.89 0.71 0.71 0.57 0.6 0.66 0.72 0.8 0.55 0.70
R 0.75 0.9 0.72 0.71 0.56 0.59 0.66 0.73 0.81 0.56 0.70
F 0.75 0.89 0.71 0.71 0.56 0.6 0.66 0.72 0.8 0.56 0.70
Table 9: Performance comparison across 10 Indic scripts on the BSTD dataset for end-to-end scene text recognition. We report Precision (P), Recall (R), and F1-score (F) for each script and the average performance. The proposed baseline model, [Uncaptioned image], significantly outperforms available open-source models and achieves competitive performance compared to commercial models. The inclusion of oracle Text Detection (TD) and Script Identification (SI) shows the potential for further improvement. The best scores for closed-source and open-source are shown in bold and underlined, respectively. The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.
4.4 End-to-End Evaluation
We now present end-to-end evaluation results on the BSTD benchmark. For this, we compare our baseline IndicPhotoOCR with the following methods:
1. Closed-Source Commercial Baselines: We evaluated two commercial systems, namely GPT-4 with vision [42] and Google OCR [41] on the BSTD benchmark [^8] We obtained their results using their best available API on October 31, 2025.. Although these comparisons are not entirely fair due to the opaque nature of their training data and model architectures, they offer a useful reference point to gage the current state of the art in Indian language scene text recognition.
We leveraged OpenAI’s GPT-4 with Vision (accessed via the gpt-4o model) to perform zero-shot scene text recognition directly. We use the following structured prompt: “Extract all text from this image. Return all words in a list format like this: [‘word1’, ‘word2’]”, with a system message defining the model’s role as: “You are an OCR system that extracts text from images.” The interaction was handled through OpenAI’s ChatCompletion.create() API, where both the instruction and the encoded image were passed as part of the message parameter. The response of the model was restricted using max tokens = 500. Although with more detailed prompt such as also locating text or chain-of-thoughts, this baseline may be improved. However, such prompt engineering is beyond the scope of this paper.
Further, we also utilized the Google Cloud Vision API to perform OCR on BSTD (Test Set).
Figure 10: A selection of results using Refer to caption. Detected bounding boxes are shown in green, with the recognized text displayed in red on a white background adjacent to each box. (Best viewed in color).
Figure 11: Failure cases in end-to-end scene text processing: The images showcase representative errors observed across the pipeline. (1) Text detection failures from TextBPN++, including false positives, missed detections, merged text, and large text misdetections. (2) Script identification errors from the ViT-based model, highlighting misclassified scripts, multi-script confusion, and challenges with degraded text. (3) Recognition failures from PARSeq, demonstrating incorrect predictions, missing or hallucinated characters, and cross-script misinterpretations. (Best viewed in color).
2. Open-Source Baselines: Currently, apart from our own, no open-source system is explicitly designed for Indian language scene text recognition. However, we include comparisons with general-purpose OCR tools such as Tesseract [37] and SuryaOCR [40], which are widely used in practical settings.
To further analyze the performance of [Uncaptioned image], we also introduce the following oracle variants: (i) IndicPhotoOCR (+Oracle TD): Uses ground-truth text detection. (ii) IndicPhotoOCR (+Oracle TD + Oracle SI): Uses ground-truth text detection and script identification. These variants help estimate the upper bound on performance in scenarios where text detection or both text detection and script identification are solved reliably.
Comparison with the aforementioned competitive baselines is presented in Table 8 and Table 9 using WRR/CRR-based metric and Precision-Recall-based metric, respectively. We observe that [Uncaptioned image] consistently outperforms all existing open-source models across all the evaluated languages. Remarkably, it also surpasses the popular commercial vision-language model GPT-4, which may not be optimized for Indian scripts. However, our approach still lags behind Google OCR in some cases, highlighting a performance gap that we aim to address in future work. Encouragingly, the variant IndicPhotoOCR (+oracle TD + SI) achieves results that are competitive, infact 30% (WRR) superior to Google OCR on an absolute scale on average, indicating the potential of our pipeline with improved text detection and script identification. In Table 9, on P/R/F metric, [Uncaptioned image] significantly surpasses other open-source baselines. The proposed baseline also not significantly behind the commercial state-of-the-art Google OCR. In fact, when combined with oracle text detection and oracle script identification, our baseline can significantly outperform Google OCR, highlighting its potential for further improvement.
A somewhat surprising observation is the noticeable performance drop for English in the end-to-end setting. Although the proposed baseline achieves a cropped word recognition accuracy of 92%, its end-to-end performance declines sharply. This drop is primarily attributable to script identification and text detection errors introduced in the full pipeline. Notably, this pattern is also reflected in commercial OCR systems such as Google’s OCR.
We also performed a comprehensive qualitative analysis of the proposed baseline. A selection of results is shown in Fig. 10. We observe that [Uncaptioned image] demonstrates strong multilingual scene text recognition across diverse real-world settings - from street signs and transportation hubs to institutional boards and cultural heritage sites. The model performs well even on slanted or perspective-distorted text (e.g., directional signs, railway nameplates), as well as on low-resolution or distant text, indicating robustness in challenging visual conditions.
On the BSTD test set, we additionally asked annotators to label each image as easy or hard based on the readability of the scene text. In the easy subset, the proposed [Uncaptioned image] achieves 60% WRR, whereas in the hard subset the performance drops to 31% across all 10 scripts. Script-wise word recognition rates are presented in Fig. 12. This clear separation between easy and hard cases highlights the sensitivity of current methods to challenging visual conditions and calls for more robust recognition models.
Figure 12: Average Word Recognition Rate (WRR) based on difficulty level.(Best viewed in color).
Figure 13: Recognition errors (words shown in red) from Google Vision, GPT-4, and Refer to caption in comparison with ground truth annotations on images from BSTD test set. (Best viewed in color).
4.5 Error Analysis
The proposed baseline is not flawless and leaves several avenues for improvement. We performed an extensive error analysis of each module of our pipeline (Section 4.4), summarized in Fig. 11, and highlight the main failure modes below.
Text Detection: Common issues in text detection include (i) false positives where background edges resemble text, (ii) missed detections in low-contrast regions, (iii) merged detections when nearby words are grouped together, and (iv) difficulty with extremely large text, suggesting the model is tuned more for small or medium size words.
Script Identification: (i) Misclassifications occur among visually similar scripts (e.g. Gujarati-Hindi, Bengali-Assamese). (ii) Multilingual overlap leads to errors when multiple scripts co-occur, as the model assumes a single script per region. (iii) Degraded or stylized text often confuses the classifier due to missing or distorted characters.
Text Recognition: Errors include (i) incorrect predictions in complex scripts like Telugu or Tamil, (ii) hallucinated or missing characters, and (iii) failures on eroded/unclear text.
These findings emphasize the need for cross-module optimization and feedback-driven refinement between detection, script identification, and recognition. In addition, these qualitative analysis, as well as the quantitative analysis in the earlier section, suggest that improving each module in the pipeline can significantly improve overall performance. These observations also motivated a visual performance comparison of IndicPhotoOCR with closed-source applications relative to the ground truth, as shown in Fig. 13.
5 [Uncaptioned image]: Installation and Usage
We open-sourced [Uncaptioned image], a toolkit for recognizing text in English and 11 Indian languages: Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Punjabi, Tamil, and Telugu. [^9] Future versions will include Urdu and Meitei. It is designed to handle the unique scripts and complex layouts of Indian languages. By releasing this toolkit publicly, we enable reproducible research, provide a common benchmark for multilingual scene text, and support the development of real-world applications across diverse Indian scripts.
We provide a user-friendly installation and usage platform, allowing our proposed baseline to be easily replicated. We firmly believe that this open-source effort will accelerate future research in Indian language scene text recognition. The installation and usage instructions are provided below.
5.1 Installation
\$ git clone https://github.com/Bhashini-IITJ/IndicPhotoOCR.git
\$ cd IndicPhotoOCR
\$ chmod +x setup.sh
\$ ./setup.sh
5.2 Usage
Detection
>>> from IndicPhotoOCR.ocr import OCR
# Initialize the OCR system
>>> ocr_system = OCR(verbose=True)
# Define input image path
>>> image_path = "path/to/image.jpg"
# Perform text detection
>>> detections = ocr_system.detect(image_path)
# Save and visualize the detection results
>>> ocr_system.visualize_detection(image_path, detections)
# Output image saved
Script Identification
>>> from IndicPhotoOCR.ocr import OCR
# Initialize the OCR system with automatic language identification model
>>> ocr_system = OCR(identifier_lang="auto")
>>> image_path = "path/to/cropped_image.jpg"
# Perform script identification
>>> recognized_lang = ocr_system.identify(image_path)
# Output recognized words
>>> print(recognized_lang)
Cropped Word Recognition
>>> from IndicPhotoOCR.ocr import OCR
>>> ocr_system = OCR(verbose=True)
>>> image_path = "path/to/cropped_image.jpg"
>>> language = "hindi"
# Perform text recognition
>>> recognized_text = ocr_system.recognise(image_path, language)
# Output recognized text
>>> print(recognized_text)
End-to-End Scene Text Recognition
>>> from IndicPhotoOCR.ocr import OCR
# Initialize the OCR system with automatic language identification
>>> ocr_system = OCR(identifier_lang="auto")
# Define input image path
>>> image_path = "path/to/image.jpg"
# Perform end-to-end OCR
>>> recognized_words = ocr_system.ocr(image_path)
# Output recognized words
>>> print(recognized_words)
Detailed documentation for usage is also available here [^10] https://bhashini-iitj.github.io/IndicPhotoOCR/.
6 Conclusion and Future Work
In this work, we introduced the Bharat Scene Text Dataset (BSTD), a large-scale benchmark specifically designed to address the gap in Indian language scene text recognition. BSTD comprises 6,582 scene images containing 1,06,478 words across 11 Indian languages and English, making it the most comprehensive dataset of its kind. Through rigorous manual annotation, we ensured high-quality polygon-level bounding boxes and text transcriptions. We also provided extensive benchmarks using state-of-the-art text detection, script identification, cropped word recognition, and end-to-end scene text recognition methods, offering a strong foundation for future research.
Our evaluations highlight the challenges posed by Indian scripts, including complex character structures, diverse vocabulary, and script variations. The results also demonstrate that existing models, primarily designed for English and Latin-based scripts, struggle with Indian languages, emphasizing the need for improved recognition techniques. Despite our extensive efforts, several challenges remain open for exploration in the future. We list them here:
Dataset Expansion: While BSTD provides a strong foundation, future extensions could include more languages (especially Urdu and Meiti) and larger-scale data to improve generalization.
Improved Recognition Models: Developing novel architectures that better capture the complexities of Indian scripts, including joint modeling of script identification and recognition, remains an open research avenue.
Benchmarking and Community Involvement: We encourage the research community to contribute to BSTD by proposing new models, sharing additional annotations, and utilizing the dataset for novel applications such as scene text translation, Indian language scene text aware captioning and question answering, and cross-lingual scene text-aware image retrieval.
We firmly believe that the BSTD and open-source toolkit [Uncaptioned image] will serve as catalyst for future advancements in the Indian language scene-text recognition and associated problems, inspiring researchers to build robust and scalable solutions for multilingual photoOCR.
Acknowledgments
This work is supported by the Ministry of Electronics and Information Technology, Govt. of India, under the NLTM-Bhashini Project.

### v2

@@
 To address these gaps, we introduce the Bharat Scene Text Dataset (BSTD) – a large-scale and comprehensive benchmark for studying Indian Language Scene Text Recognition.
-It comprises more than 100K words that span 11 Indian languages and English, sourced from over 6,500 scene images captured across various linguistic regions of India.
+It comprises more than 100K words that span 11 Indian languages and English, sourced from over 6,582 scene images captured across various linguistic regions of India.
 The dataset is meticulously annotated and supports multiple scene text tasks, including: (i) Scene Text Detection, (ii) Script Identification, (iii) Cropped Word Recognition, and (iv) End-to-End Scene Text Recognition.
@@
 Our results highlight the challenges and opportunities in Indian language scene text recognition.
-We believe that this dataset represents a significant step toward advancing research in this domain.
+We believe that our dataset represents a significant step toward advancing research in this domain.
 All our models and data are open source.
@@
 This figure showcases typical street scenes from northern, eastern, western, and southern parts of India, highlighting the country’s vast linguistic diversity.
-With 11 languages (five of them: Hindi, English, Bengali Gujarati and Tamil are shown here) including English commonly featured on signboards, scene text understanding in India presents unique challenges.
-While significant progress has been made in English Scene Text Recognition, open-source comprehensive effort for Indian language scene text understanding, including large public datasets and comprehensive models, are still limited.
-Our work seeks to address this gap and advance the field.
+With 11 languages including English commonly featured on signboards, scene text understanding in India presents unique challenges.
+Scene Text from five languages, namely Hindi, English, Bengali Gujarati and Tamil are shown here.
+While significant progress has been made in English Scene Text Recognition, open-source comprehensive effort for Indian language scene text understanding, including large public datasets and open-source models, are still limited.
+Our work seeks to address this gap toward advancing the field.
 1 Introduction
-Scene Text Recognition, especially for English, has advanced significantly in the last decade, thanks to a series of efforts on dataset creation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and model development [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
+Scene Text Recognition, especially for English, has advanced significantly in the last decade, thanks to a series of efforts on dataset creation [ICDAR2003, SVT, mishra2012scene, MSRA-TD500, ICDAR2013, ICDAR2015, CUTE80, Total-Text, SVT-P, MLT-19] and model development [cnn_detector_2014, CRNN2016, jaderberg2016reading, east, textspotter2018, craft, ASTER2019, ABCNet2020, dbnet, visionLAN2021, ABINet2021, parseq, ABINet++2023, textbpn++, clip4str].
 In contrast, scene text recognition for Indian languages has been extremely underexplored and possibly more challenging (refer to Fig.
@@
 Note that Indian languages are used by more than 1.4 billion people, which is approximately 18% of the global population.
-Despite its usage and popularity, we observe that scene text recognition in Indian languages lacks a comprehensive benchmark and open-source models, except for some early efforts [26, 27, 28, 29, 30, 31, 32, 33] that laid important groundwork, but did not aim for a comprehensive coverage of tasks or languages.
+Despite its usage and popularity, we observe that scene text recognition in Indian languages lacks a comprehensive benchmark and open-source models, except for some early efforts [recognition-ISI-umapada2010, IIIT_ILST, detection-ISI-Umapada2017, kakwani-etal-2020-indicnlpsuite, IIIT-TL-STR, IIIT-indicSTR12, VijayanCDK25] that laid important groundwork, but did not aim for a comprehensive coverage of tasks or languages.
 We aim to fill this gap through our work.
@@
 The BSTD comprises 6.5K scene images that feature more than 100K words in 11 Indian languages, along with English.
-These images, sourced from Wikimedia [^2] https://www.wikimedia.org/, span various signboards and accidental background text from different Indian cities and towns, ensuring that the dataset captures the linguistic diversity of India.
-We obtain a polygon-level bounding box for each visible text in the image and the text it contains using rigorous manual annotation.
-The dataset is publicly available for download here [^3] https://github.com/Bhashini-IITJ/BharatSceneTextDataset.
+These images that are sourced from Wikimedia [^2] https://www.wikimedia.org/, span various signboards and accidental background text from different Indian cities and towns, ensuring that the dataset captures the linguistic diversity of India.
+For each visible text instance in an image, we manually annotate a polygonal bounding region and provide its corresponding transcription.
+The dataset is publicly available for download [^3] https://github.com/Bhashini-IITJ/BharatSceneTextDataset.
 The BSTD can be used for the following scene text tasks: (i) scene text detection, (ii) script identification, (iii) cropped word recognition, and (iv) end-to-end scene text recognition.
-In this work, we benchmark the BSTD using our new pipeline, [Uncaptioned image], which integrates state-of-the-art architectures across its modules.
+In this work, we benchmark the BSTD using our new pipeline, [Uncaptioned image], which integrates state-of-the-art architectures in its modules.
 We firmly believe that these benchmarks will help accelerate advancements in Indian language scene text recognition.
+Table 1: Comparison of the Bharat Scene Text Dataset (BSTD) with existing scene text datasets.
+BSTD supports all three core tasks - text detection, script identification, and recognition (and, thereby end-to-end evaluation) across 11 Indian languages along with English.
+It is the only publicly available dataset for Indian languages that enables comprehensive end-to-end scene text understanding.
+: MLT-17 and MLT-19 are not focused on Indian languages, they only include Hindi and Bengali.
+\$: IIIT-IndicSTR12 and #: IIIT-STR [GunnaSJ22] are primarily annotated for cropped word recognition; consequently, neither can be utilized to evaluate end-to-end scene text recognition as ours.
 Dataset # Images # Words # Languages Tasks Supported
 Detection Script Identification Recognition
-ICDAR 2003 [1] 509 2,268 1
-SVT [2] 350 725 1
-IIIT-5K [3] - 5,000 1
-ICDAR 2013 [5] 462 5,003 1
-ICDAR 2015 [6] 1,500 6,545 1
-CUTE80 [7] 80 288 1
-Total-Text [8] 1,555 11,459 1
-IIIT-TL-STR [30] 440 3,035 2
-IIIT-ILST [27] - 3,168 3
-MLT-17 [34] 18,000 96,000 9
-MLT-19 [10] 20,000 191,000 10
-IIIT-indicSTR12∗∗ [31] - 27,000 12
-BSTD (This Work) 6,582 1,06,478 12
-Table 1: Comparison of the Bharat Scene Text Dataset (BSTD) with existing scene text datasets.
-BSTD supports all three core tasks—text detection, script identification, and recognition (and, thereby end-to-end evaluation) across 12 Indian languages, including English.
-It is the only publicly available dataset for Indian languages that enables comprehensive end-to-end scene text understanding.
-: MLT-17 and MLT-19 are not focused on Indian languages, they only include Hindi and Bengali.
-**: IIIT-IndicSTR12 is primarily annotated for cropped word recognition and can not be studied to evaluate end-to-end scene text recognition as ours.
+ICDAR 2003 [ICDAR2003] 509 2,268 1
+SVT [SVT] 350 725 1
+IIIT-5K [mishra2012scene] - 5,000 1
+ICDAR 2013 [ICDAR2013] 462 5,003 1
+ICDAR 2015 [ICDAR2015] 1,500 6,545 1
+CUTE80 [CUTE80] 80 288 1
+Total-Text [Total-Text] 1,555 11,459 1
+IIIT-TL-STR [IIIT-TL-STR] 440 3,035 2
+IIIT-ILST [IIIT_ILST] - 3,168 3
+MLT-17 [MLT-17] 18,000 96,000 9
+MLT-19 [MLT-19] 20,000 191,000 10
+IIIT-indicSTR12\$ [IIIT-indicSTR12] - 27,000 12
+IIIT-STR# [GunnaSJ22] - 12,057 6
+BSTD (This Work) 6,582 1,26,292 12
 Figure 2: Language map of India (Source: Wikimedia).
@@
   1. 1.
-    We introduce the Bharat Scene Text Dataset (BSTD) – a large-scale benchmark for Indian language scene text recognition, consisting of 6,582 scene images featuring 1,06,478 words in 11 Indian languages and English.
-Each image is manually annotated with polygon-level bounding boxes and corresponding text and script, ensuring high-quality data for research and applications.
+    We introduce the Bharat Scene Text Dataset (BSTD) – a large-scale benchmark for Indian language scene text recognition, consisting of 6,582 scene images featuring 1,26,292 words in 11 Indian languages and English.
+Each image is manually annotated with polygon-level bounding boxes and corresponding transcription and script, ensuring high-quality data for research and applications.
 (Section 2)
   2. 2.
+    We present a strong baseline and release corresponding open-source toolkit, namely [Uncaptioned image], designed to detect, identify, and recognize text in English and 11 Indian languages.
+The released toolkit is easy to install and replicate, making it accessible to researchers and developers.
+(Section 3 and Appendix-A)
+  3. 3.
     We evaluated BSTD using state-of-the-art techniques, including off-the-shelf methods and fine-tuned models on synthetic and/or BSTD train data (as applicable).
-Detailed benchmark results are provided for all key tasks, i.e., text detection, script identification, cropped word recognition, and end-to-end scene text recognition.
+Detailed benchmark results including failure case analysis are provided for all key tasks, i.e., text detection, script identification, cropped word recognition, and end-to-end scene text recognition.
 These evaluations establish a strong baseline for future research and advancements in multilingual scene text recognition, particularly for Indian languages.
 (Section 4)
-  3. 3.
-    We introduce an open source toolkit, namely [Uncaptioned image], designed to detect, identify, and recognize text in English and 11 Indian languages.
-The toolkit is easy to install and replicate, making it accessible to researchers and developers.
-(Section 5)
+Figure 3: Pipeline for constructing a multilingual scene-text dataset from Wikimedia Commons images.
+The process begins with query construction using combinations of Indian city names and place-related keywords to retrieve relevant images via the Wikimedia Commons API.
+A two-stage filtering process using DBnet [dbnet] (pre-trained on Indic Synthetic Data) removes non-textual images and manually filters relevant ones.
+Scene-text detection annotations are generated using DBnet and refined through manual corrections.
+Finally, cropped word images are recognized using the PARSeq [parseq] model (trained on 11 Indic languages using synthetic data), with recognition errors corrected and script language tags added to produce the final annotated dataset.
 2 Bharat Scene Text Dataset
@@
 In addition, the identification of scripts becomes extremely important.
-Although there has been some effort in collecting and benchmarking Indian language scene text understanding tasks, it falls short compared to efforts for scene text understanding in English [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
-Please refer to Table 1 for a detailed comparison of the Bharat Scene Text Dataset compared to other related scene text datasets.
+Although there has been some effort in collecting and benchmarking Indian language scene text understanding, such as [IIIT_ILST, IIIT-TL-STR, IIIT-indicSTR12, GunnaSJ22], it falls short compared to efforts for scene text understanding in English [ICDAR2003, SVT, mishra2012scene, MSRA-TD500, ICDAR2013, ICDAR2015, CUTE80, Total-Text, SVT-P, MLT-19].
 To address this gap, we present the Bharat Scene Text Dataset, a novel and comprehensive dataset specifically curated to represent the linguistic diversity of India.
 By using the term “Bharat”, we not only acknowledge the nation’s deep-rooted cultural identity but also highlight our focus on supporting scene text understanding across 11 major Indian languages.
-This dataset contains 6,582 scene images and 1,06,478 word images in total.
+This dataset contains 6,582 scene images and 1,26,292 word images in total.
 This dataset and associated annotations can be downloaded from our project website [^4] https://vl2g.github.io/projects/IndicPhotoOCR/.
-In this section, we give details of dataset curation and annotation, supported tasks, train-test splits followed by dataset analysis.
+Please refer to Table 1 for a detailed comparison of the Bharat Scene Text Dataset compared to other related scene text datasets.
+In the next section, we give details of dataset curation and annotation, supported tasks, train-test splits followed by dataset analysis.
 2.1 Dataset Curation and Annotation
 We illustrate the pipeline for constructing the Bharat Scene Text Dataset from Wikimedia Commons in Fig. 3. The dataset is curated and annotated in the following three stages:
-Figure 3: Pipeline for constructing a multilingual scene-text dataset from Wikimedia Commons images.
-The process begins with query construction using combinations of Indian city names and place-related keywords to retrieve relevant images via the Wikimedia Commons API.
-A two-stage filtering process using DBnet [19] (pre-trained on Indic Synthetic Data) removes non-textual images and manually filters relevant ones.
-Scene-text detection annotations are generated using DBnet and refined through manual corrections.
-Finally, cropped word images are recognized using the PARSeq [22] model (trained on 11 Indic languages using synthetic data), with recognition errors corrected and script language tags added to produce the final annotated dataset.
 2.1.1 Stage-1: Image collection
-We compile a comprehensive list of cities and towns in various regions of India where 11 selected languages are spoken predominantly.
+We compile a comprehensive list of cities and towns across various regions of India where 11 selected languages are spoken predominantly.
 Furthermore, we curate a catalog of public places, such as ‘bus stations’, ‘banks’, ‘schools’, ‘railway stations’, ‘hospitals’, ‘ATMs’, ‘metro stations’, ‘street signs’, ‘airport’, ‘shop entrance’, ‘billboards’, ’press conference’, etc.
 By combining the city and place names as search queries, e.g.
-‘Delhi Railway Station’, we retrieved images along with their metadata from the Wikimedia Commons API [^5] https://commons.wikimedia.org/w/api.php.
-It should be noted that not all retrieved images contain scene text at this stage, which requires subsequent filtering.
-At this stage, we obtain 50K images.
+‘Delhi Railway Station’, ‘Kolkata Bookshop’, we retrieved images along with their metadata from the Wikimedia Commons API [^5] https://commons.wikimedia.org/w/api.php.
+It should be noted that not all retrieved images contain scene text, necessitating a subsequent filtering process.
+At this stage, we obtain $\approx$ 50K images.
 Figure 4: An example image from BSTD along with the corresponding JSON annotation.
 2.1.2 Stage-2: Image filtering
-We filter the images harvested in the previous step using the pre-trained scene text detection method [19].
+We filter the images harvested in the previous step using the pre-trained scene text detection method [dbnet].
 To this end, we run this detector and obtain text even with low confidence.
 We retain images containing at least one word, while others were discarded.
-In addition, we ask human annotators to review the filtered set to eliminate false positives.
+Nevertheless, 53 images in the final dataset were found to contain no annotatable scene text.
+We deliberately retained this small fraction of non-text images to preserve the natural variability of real-world data collection pipelines, where some captured scenes may not ultimately contain visible text.
 The final data set comprises 6,582 images that contain scene text in 11 Indian languages and English.
@@
 We obtain bounding box annotation using a fine-tuned DBnet model on synthetic Indian language data [^6] Refer to Section-3.3.1 for details of synthetic data generation.
-Since these automatically obtained bounding boxes are not always perfect, they are refined by our hired annotators using an in-house tool to edit, draw, or delete polygons around the scene text words.
-Once the bounding boxes were obtained, they were assigned to the annotators and assigned to label each box with the script it contains.
+Since these automatically obtained bounding boxes are not always perfect, they were refined by our hired annotators using an internal tool to edit, draw, or delete polygons around the scene text words.
+Once the bounding boxes were obtained, they were assigned to the annotators and tasked to label each box with the script it contains.
 These annotators were well-versed in script identification and were also informed about the geographical origin of each image.
-With this context, they were able to accurately label each bounding box with the corresponding script.
-We then use PARSeq [22], fine-tuned on synthetic data for Indian languages, to read the text within each bounding box.
+With this context, they were able to accurately label each bounding box with theier corresponding script.
+We then use PARSeq [parseq], fine-tuned on synthetic data for Indian languages, to read the text within each bounding box.
 The bounding boxes, along with the recognized text, are then provided to language expert annotators.
 These annotators review the text and correct any recognition errors or retain the text as is if it is accurate.
+In addition, for every image, we have included annotation on the lighting (captured during the day or night) and environment (captured outdoor or indoor) conditions.
+Fig. 4 shows the fine grained nature of the annotations done in this work.
 2.2 Annotator Details
@@
 The entire process–pseudo-annotation, manual correction, and quality verification took 10 months.
-Fig 4 shows the fine grained nature of the annotations done in this work.
 2.3 Supported Tasks and Evaluation Protocols
@@
 We provide polygon-level tight bounding boxes for each word image in the dataset.
-The popular metrics such as Precision, Recall and F1-score using TedEval evaluation [35] can be used to assess the performance of the method.
+The popular metrics such as Precision, Recall and F1-score using TedEval evaluation [tedeval] can be used to assess the performance of the scene text detectors.
   2. 2.
@@
     End-to-End Scene Text Recognition: This task is the closest to the natural setting.
-Here, given a scene image, the task is to recognize all the instances of text.
+Here, given a scene image, the task is to recognize all instances of text.
 Solving end-to-end Scene Text Recognition for Indian languages, often requires a pipeline of scene text detection, script identification followed by cropped word recognition.
@@
 Note that precision is defined as the number of correctly recognized words over the total number of recognized words, while recall is defined as the number of correctly recognized words over the total number of words in the ground truth.
-Figure 5: Box plot highlighting the word length range across all the 12 languages.
+Table 2: Summary of bounding box annotation in the BSTD, detailing the total, training, and test set distributions
+Set Images Bounding Box Annotations
+Total 6,582 1,26,292
+Train 5,263 94,128
+Test 1,319 32,164
 We randomly divided the 6,582 images into an 80-20% train-test split.
@@
 The split comprises 1,800 training images and 478 testing images per language.
-Set Images Bounding Box Annotations
-Total 6,582 1,26,292
-Train 5,263 94,128
-Test 1,319 32,164
-Table 2: Summary of image and detection annotation counts in the BSTD dataset, detailing the total, training, and test set distributions
+Figure 5: Box plot highlighting the word length range across all the 12 languages.
+Table 3: Cropped word image data split for the BSTD, showing the number of annotated instances, train images, and test images across multiple languages.
+Here, Others correspond to those words which are not from the listed languages here (e.g., Meitei and Urdu).
+We plan to include their annotation in subsequent versions of our dataset.
+For scene-level language distribution statistics, please refer to Table 4.
 Language #Word images Train Test
@@
 Total 1,26,292 77,726 28,752
-Table 3: Cropped word image data split for the BSTD dataset, showing the number of annotated instances, train images, and test images across multiple languages.
-Here, Others correspond to those words which are not from the listed languages here (e.g., Meitei and Urdu).
-We plan to include their annotation in subsequent versions of our dataset.
-Figure 6: Distribution of the bounding box size in our dataset, in terms of pixels.
+Figure 6: Distribution of word-level bounding box areas in the BSTD, plotted on a base-10 logarithmic scale.
+Table 4: Scene-level language composition of the BSTD.
+The distribution highlights the multilingual nature of the dataset, with a significant majority of images containing English co-occurring with one or more Indian languages.
+Language Composition #Images
+English only 53
+Indian languages only 1,958
+English + one or more Indian languages 4,518
+No annotatable scene text 53
+Total 6,582
 2.4 Dataset Analysis
@@
 We observe that it has a good coverage of all the major Indian languages with minimum 2,542 word images for Tamil.
-This makes the dataset linguistically rich.
+Further, we performed an image-level analysis and observed that, out of the 6,582 total scene images in the dataset, only 53 contain exclusively English text, 1,958 contain only non-English text (i.e., Indian languages), and 4,518 contain English co-occurring with one or more Indian languages.
+This makes the dataset linguistically rich (refer to Table 4).
 We further analyze the average word length for each language in our dataset in Fig.
-5).
+5.
 The plot reveals that Malayalam, Tamil, and Telugu contain longer words on average, whereas English and Hindi have shorter word lengths.
@@
 6.
-It indicates that most text regions (59.4%) are small (1001–10,000 pixels), while only 2.6% of text instances fall into the largest category (above 100,000 pixels).
-These findings highlight the diverse textual characteristics within BSTD and it contains a mix of small and large text regions within images making it challenging for detection.
-Furthermore, the high percentage of small text regions suggests that many images contain incidental text rather than just focused text, which adds to the dataset’s difficulty implying the real word scenario.
-Figure 7: Word cloud for recognition annotations in our BSTD dataset: Highlighting the diversity of vocabulary across 12 languages.
-The tiles are arranged row-wise from left to right as follows: the first row contains Hindi, Assamese, and Gujarati; the second row shows English, Malayalam, and Kannada; the third row corresponds to Odia, Panjabi, and Tamil; and the fourth row includes Telugu, Marathi, and Bengali.
-Figure 8: A selection of challenging cropped words from our dataset.
-Few key challenges include eroded text (c, d, k), handwritten text (f), occluded text (e, h, l), ambiguous text (a, b, i), text in different orientations (g, j).
-Finally, to illustrate diversity in the dataset, we show word clouds for all 12 languages in our data set in Fig.
-7.
-As can be seen, the most frequent words (represented by the largest font size) across all languages point to text commonly found on public signs, government buildings, and commercial establishments.
-In addition, the vocabulary is dominated by nouns that are staples of scene text recognition tasks.
-These are words that one would expect to see “in the wild”.
-Some of the cropped word samples from BSTD samples are shown in Fig.
-8.
+It indicates that bounding box sizes range from very small text region, e.g., minimum area = 32 pixels to very-large text region, e.g., maximum area $\approx$ $2.49\times 10^{6}$ pixels.
+The distribution closely approximates a log-normal curve that peaks near the median area of 2,889 pixels.
+These findings highlight that the BSTD contains a mix of small and large text regions within images, making it challenging for detection.
+Furthermore, the high proportion of small text regions indicates the prevalence of incidental text, as opposed to only focused text, thereby making the dataset more challenging and representative of real-world scenarios.
+An analysis of the lighting and environmental conditions of the scene images revealed that approximately 89% of the images were captured during daylight, whereas the remaining 11% were taken at night.
+Further, the environment annotations indicate that 91% of the images correspond to outdoor scenes, while only 9% were captured indoors.
 In short, the analysis of the Bharat Scene Text Dataset suggests that it has diverse linguistic and geographical coverage, with a clear focus on text from public, commercial and transport-related signage, and has a mix of focused and incidental text, making it a valuable resource for developing robust real-world text recognition models for Indian languages.
 Further, while widely spoken languages such as Hindi, English, Bengali, and Marathi are well represented in the dataset, low-resource languages including Assamese, Gujarati, Kannada, Malayalam, Odia, Punjabi, Tamil, and Telugu also have a substantial presence, with at least 2,542 word images for Tamil (see Table 3).
-This shows that the dataset is not only strong in high-resource languages but also sufficiently representative of low-resource languages, making it suitable for comprehensive multilingual text recognition research.
+This shows that the dataset is not only strong in high-resource languages, but also sufficiently representative of low-resource languages, making it suitable for comprehensive multilingual text recognition research.
 3 [Uncaptioned image]: Proposed Baseline
 Scene text recognition in a multilingual environment typically involves three key modules: (i) Scene Text Detection, (ii) Script Identification, and (iii) Cropped Word Recognition.
-In the absence of a dedicated architecture for Indian language scene text recognition in the literature, we develop a strong baseline by adapting or fine-tuning (as needed) state-of-the-art architectures across various modules.
+In the absence of a dedicated architecture for Indian language scene text recognition in the literature, we develop a strong baseline by adapting or fine-tuning (as needed) state-of-the-art architectures across these modules.
 We describe each of these modules below and collectively refer to this pipeline as [Uncaptioned image].
@@
 Hence, we directly incorporate them into our pipeline without language-specific modifications.
-In particular, we adopt TextBPN++ [24] for scene text detection, as it demonstrated superior performance compared to other detectors in our evaluations.
+In particular, we adopt TextBPN++ [textbpn++] for scene text detection, as it demonstrated superior performance compared to other detectors in our evaluations.
 Detailed results for this stage are presented in the experimental section 4.1.
@@
 We fine-tuned the model, including both the transformer backbone and the classification head, using the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a batch size of 16 using BSTD-train set.
-This ViT-based approach consistently outperformed other configurations, demonstrating strong generalization in both script classification settings of the 3-way and 12-way scripts.
-Language Train Size Test Size Vocabulary Size #Fonts
+This ViT-based approach consistently outperformed other configurations, demonstrating strong generalization in both 3-way and 12-way script classification settings.
+Table 5: Statistics for Synthetic Dataset generated using [synthtext] for 11 Indian languages.
+Language Train Size Test Size Vocab Size #Fonts
 Assamese 8,092,205 899,134 636,312 44
@@
 Telugu 3,485,570 387,286 5,705,371 42
-Table 4: Statistics for Synthetic Dataset generated using [36] for 11 Indian languages.
 3.3 Cropped Word Recognition
 Once text instances are detected and their scripts identified, the next step is text recognition.
-Unlike detection, this task relies heavily on language-specific modeling and requires large-scale training data.
-Off-the-shelf models for English text recognition cannot be directly applied as they are not designed for this task.
-Instead, we fine-tune a successful scene text recognition architecture (namely PARSeq [22]) for the 11 Indian languages in our study.
+In contrast to detection, this task depends strongly on language-specific modeling and demands large-scale training data.
+Existing off-the-shelf English scene text recognition models are not directly applicable, as they are not designed for Indian scripts.
+To address this, we fine-tune a successful scene text recognition architecture, PARSeq [parseq], for the 11 Indian languages included in our study.
 A key challenge in this setting is the scarcity of real-world scene text data for low-resource languages, in contrast to English, where recognizers are trained on millions of samples.
@@
 This lack of data led us to adopt an existing method for generating synthetic scene text images.
-We used SynthText [36] to generate large-scale data, with statistics summarized in Table 4.
-SynthText [36] provided 8000 background images, which we leveraged as the background image for our synthetic data.
+We used SynthText [synthtext] to generate large-scale data, with statistics summarized in Table 5.
+SynthText [synthtext] provided 8000 background images, which we leveraged as the background image for our synthetic data.
 To generate scene text in different Indian languages, we needed appropriate vocabulary and fonts for each language.
-We sourced vocabulary from the AI4Bharat-IndicNLP Dataset [^7] https://github.com/AI4Bharat/indicnlp_corpus introduced in the IndicNLPSuite [29].
+We sourced vocabulary from the AI4Bharat-IndicNLP Dataset [^7] https://github.com/AI4Bharat/indicnlp_corpus introduced in the IndicNLPSuite [kakwani-etal-2020-indicnlpsuite].
 SynthText also accounts for local 3D geometry, enabling varied orientation of text placement within the scene.
@@
 In this section, we provide a detailed evaluation of the proposed baseline, namely [Uncaptioned image], and its individual modules on the newly introduced Bharat Scene Text Dataset.
-We compare it with several open-source OCR systems, including Tesseract [37], PaddleOCR [38], EasyOCR [39], and SuryaOCR [40].
-For end-to-end evaluation, we also compare it with commercial closed-source systems such as Google OCR [41] and GPT-4 [42].
+We compare it with several open-source OCR systems, including Tesseract [tesseract], PaddleOCR [cui2025paddleocr30technicalreport], EasyOCR [jaidedai_easyocr_2020], and SuryaOCR [paruchuri2025surya].
+For end-to-end evaluation, we also compare it with commercial closed-source systems such as Google OCR [googlecloud_visionocr] and GPT-4 [gpt_4v].
 In addition, we include appropriate competitive approaches to evaluate individual modules.
@@
 They also demonstrate the effectiveness of [Uncaptioned image] in a challenging setting.
-Please note that all the results are reported on the Bharat Scene Text Dataset (test set).
+Please note that all the quantitative results are reported on the Bharat Scene Text Dataset (test set).
+Table 6: Scene Text Detection results on BSTD-Test Set using techniques prevalent in English scene text datasets.
+Method P R F1
+EAST [east] 0.55 0.10 0.17
+CRAFT [craft] 0.51 0.22 0.19
+DBNet [dbnet] 0.71 0.50 0.59
+Hi-SAM [Hi-SAM] 0.75 0.19 0.46
+TextBPN++ [textbpn++] 0.75 0.78 0.77
 4.1 Results on Scene-text detection
-Our proposed baseline, [Uncaptioned image], employs TextBPN++ [24] for scene text detection.
-We compare its performance with several widely used scene text detection methods, including EAST [14], CRAFT [16], DBNet [19], and Hi-SAM [43].
+Our proposed baseline, [Uncaptioned image], employs TextBPN++ [textbpn++] for scene text detection.
+We compare its performance with several widely used scene text detection methods, including EAST [east], CRAFT [craft], DBNet [dbnet], and Hi-SAM [Hi-SAM].
 The comparison results are summarized in Table 6.
 We observe that TextBPN++ significantly outperforms the other detectors, achieving higher F1-score and recall.
-Although Hi-SAM is also equally precise, it misses out many in detecting many Indian language words.
-Please note that we use TedEval [35] evaluation for reporting these numbers.
-Method P R F1
-EAST [14] 0.55 0.10 0.17
-CRAFT [16] 0.51 0.22 0.19
-DBNet [19] 0.71 0.50 0.59
-Hi-SAM [43] 0.75 0.19 0.46
-TextBPN++ [24] 0.75 0.78 0.77
-Table 5: Scene Text Detection results on BSTD-Test Set using techniques prevalent in English scene text datasets.
-4.2 Results on Script Identification
-For script identification, we evaluate performance under two settings: (i) 3-way classification, where the model distinguishes among a regional language, Hindi and English.
-This setup reflects a common real-world scenario in India, where signboards often include all three languages.
-(ii) 12-way classification, where the model selects from the English and 11 Indian languages.
-We use the following baselines: (i) AlexNet [44]: a popular CNN-based Imagenet-pretrained image classification model finetuned on trainset of our script identification data, (ii) CRNN [45]: a CNN-based scene text recognition model with the ability to encode sequence information, adapted for script identification, followed by state-of-the-art transformer architectures for image classification; (iii) CLIP [46] and (iv) ViT [47], respectively.
-All the results reported in Table 6 correspond to the models trained on real images.
+Although Hi-SAM is also equally precise, it misses out detecting many Indian language words.
+Please note that we use TedEval [tedeval] evaluation in our experiments.
+Table 7: Script identification accuracy (%) in the 3-way classification setting: Hindi vs English vs one regional Indian language shown in the language column.
+Each row corresponds to a different regional Indian language, while Hindi and English remain fixed across all rows.
 Trained on BSTD
-Language AlexNet [44] CRNN [12] CLIP [46] Proposed
-Assamese 53.1 72 75.6 92.2
-Bengali 54 75.7 77.5 93.3
+Language AlexNet CRNN CLIP Proposed
+Assamese 53.1 72.0 75.6 92.2
+Bengali 54.0 75.7 77.5 93.3
 Kannada 54.7 73.8 82.7 94.5
@@
 Punjabi 51.3 81.1 79.7 92.3
-Table 6: Script Identification Accuracy in 3-way classification setting on Hindi vs English vs third regional Indian language shown in the language column.
-Figure 9: Normalized confusion matrix illustrating the performance of script identification model across 12 Indian. (Best viewed in color).
+4.2 Results on Script Identification
+For script identification, we evaluate performance under two settings: (i) 3-way classification, where the model distinguishes among a regional language, Hindi and English.
+This setup reflects a common real-world scenario in India, where signboards often include all three languages.
+(ii) 12-way classification, where the model selects from the English and 11 Indian languages.
+We use the following baselines: (i) AlexNet [alexnet]: finetuned on trainset of our script identification data, (ii) CRNN [CRNN2016]: a CNN-based scene text recognition model with the ability to encode sequence information, adapted for script identification, followed by state-of-the-art transformer architectures for image classification; (iii) CLIP [clip] and (iv) ViT [vit], respectively.
+Figure 7: Normalized confusion matrix illustrating the performance of script identification in 12-way classification setting. (Best viewed in color).
 For the 3-way classification, the results are reported in Table 7.
@@
 The corresponding normalized confusion matrix is presented in Fig.
-9.
+7.
 While the overall results are satisfactory, the model frequently confuses Assamese with Bengali and Hindi with Marathi.
@@
 Our end-to-end evaluation suggests that further improving script identification can significantly boost overall scene text recognition performance.
+Table 8: Baseline Results on Cropped Word Scene Text Recognition for 12 languages on BSTD. Please note that PARSeq is used in our proposed baseline – [Uncaptioned image].
+Model Train Data Assamese Bengali English Gujarati Hindi Kannada Malayalam Marathi Odia Punjabi Tamil Telugu Average
+PARSeq [parseq] Synth 0.47 0.44 0.92 0.32 0.43 0.47 0.52 0.44 0.42 0.4 0.37 0.44 0.47
+BSTD Train 0.79 0.82 0.92 0.61 0.71 0.6 0.58 0.86 0.72 0.75 0.8 0.56 0.73
+CRNN [CRNN2016] AI4Bharat NA NA 0.27 NA 0.13 NA 0.06 NA NA 0.27 0.16 0.05 0.16
+Off-the-shelf open source models
+Tesseract [tesseract] - 0.15 0.18 0.33 0.12 0.2 0.12 0.03 0.21 0.11 0.19 0.13 0.09 0.15
+PaddleOCR [cui2025paddleocr30technicalreport] - NA NA 0.61 NA 0.31 0.03 NA 0.36 NA NA 0.28 0.15 0.29
+EasyOCR [jaidedai_easyocr_2020] - 0.12 0.24 0.29 NA 0.18 0.06 NA 0.25 NA NA NA 0.11 0.18
+Figure 8: A selection of challenging cropped words from our dataset along with ground truth and recognized results by various methods including Refer to caption.
+Key challenges include out-of-focus text (a, b, i), varied fonts (c), eroded text (d, k), handwritten text (f), occluded text (e, h, l), and skewed text (g, j).
+Cells marked with “-” indicate an empty recognition, while “NA” denotes that specific language support was unavailable for that model.
+Here, OCR output text in red color shows that the word is incorrectly recognized.
+(Best viewed in color).
+Table 9: Comparative evaluation for end-to-end scene text recognition on BSTD using WRR and CRR.
+We present baseline results under two scenarios: (i) when text detection is provided as ground truth (denoted as oracle TD) and (ii) when both text detection and script identification are given as ground truth (denoted as TD+SI).
+“ / ” indicates whether the tool is open-source ( ) or closed-source ( ).
+Bold and underlined number shows best performing results in non-oracle setting for closed and open-source models respectively.
+The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.
+WRR on BSTD-Test
+Method / Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
+ChatGPT [gpt_4v] 0.08 0.36 0.09 0.2 0.03 0.05 0.04 0.12 0.21 0.1 0.13
+Google OCR [googlecloud_visionocr] 0.43 0.37 0.32 0.55 0.24 0.37 0.43 0.46 0.52 0.42 0.41
+Tesseract [tesseract] 0.04 0.02 0.02 0.03 0 0 0.02 0.02 0 0.02 0.02
+SuryaOCR [paruchuri2025surya] 0.18 0.15 0.09 0.19 0.07 0.12 0.15 0.21 0.24 0.05 0.14
+[Uncaptioned image] 0.45 0.37 0.3 0.39 0.29 0.31 0.38 0.41 0.42 0.31 0.36
+[Uncaptioned image] (+oracle TD) 0.75 0.74 0.54 0.63 0.49 0.54 0.68 0.7 0.69 0.49 0.62
+[Uncaptioned image] (+oracle TD+SI) 0.8 0.92 0.61 0.74 0.6 0.58 0.72 0.75 0.8 0.56 0.71
+CRR on BSTD-Test
+Method / Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
+ChatGPT [gpt_4v] 0.13 0.46 0.15 0.28 0.1 0.14 0.11 0.2 0.31 0.21 0.21
+Google OCR [googlecloud_visionocr] 0.57 0.42 0.36 0.66 0.45 0.55 0.57 0.59 0.69 0.6 0.55
+Tesseract [tesseract] 0.12 0.07 0.06 0.09 0.05 0.09 0.09 0.09 0.04 0.06 0.08
+SuryaOCR [paruchuri2025surya] 0.29 0.24 0.15 0.33 0.19 0.26 0.29 0.31 0.45 0.26 0.28
+[Uncaptioned image] 0.59 0.47 0.43 0.54 0.48 0.55 0.59 0.53 0.62 0.57 0.54
+[Uncaptioned image] (+oracle TD) 0.87 0.78 0.64 0.74 0.68 0.8 0.84 0.84 0.78 0.73 0.77
+[Uncaptioned image] (+oracle TD+SI) 0.93 0.97 0.75 0.89 0.86 0.86 0.89 0.9 0.94 0.83 0.88
 4.3 Results on Cropped Word Recognition
 For the cropped-word recognition baselines, there are no models specifically designed for Indian languages.
-As a result, we must rely on English scene-text recognition approaches such as [25, 48, 22, 45].
+Therefore, we could adopt English scene-text recognition approaches such as [clip4str, clipter, parseq, CRNN2016].
 However, adopting all these methods still demands substantial computational resources, especially when running recognition models across 12 languages.
-Therefore, we only choose PARSeq [22], one of the recent and effective recognition models, as our baseline.
-In addition, we compare the PARSeq-based baseline with the already available, CRNN [12] by AI4Bharat [49] and off-the-shelf open-source models, namely Tesseract [37], PaddleOCR [38], and EasyOCR [39].
+Therefore, we only choose PARSeq [parseq], one of the recent and effective recognition models, as our baseline.
+In addition, we also use available, CRNN [CRNN2016] by AI4Bharat [gokulkarthik_indianstr_2023] and off-the-shelf open-source models, namely Tesseract [tesseract],PaddleOCR [cui2025paddleocr30technicalreport], and EasyOCR [jaidedai_easyocr_2020].
 PARSeq is trained on a combination of synthetic data and the BSTD dataset.
-Our baseline for cropped word recognition is based on the existing text recognition framework PARSeq [22].
+Our baseline for cropped word recognition is based on the existing text recognition framework PARSeq [parseq].
 We report results under two settings: (i) training PARSeq from scratch using synthetic data, and (ii) fine-tuning it on the BSTD training set.
@@
 This highlights the critical role of real-world data in scene text recognition.
-Also, as can be seen, Word Recognition Rates for cropped English scene text are relatively high (around 92%).
+Also, as can be seen, Word Recognition Rates for cropped English scene text are relatively high (around 92%) [^8] Unless otherwise specified, the English recognition model is trained on both uppercase and lowercase characters; however, the decoder normalizes predictions to lowercase, and evaluation is performed in a case-insensitive manner.
+We additionally conducted an experiment without this relaxation, i.e., using case-sensitive decoding and evaluation.
+Under this stricter setting, the model achieves a WRR of 84%, compared to 92% under case-insensitive decoding..
 However, performance drops noticeably for other languages: Marathi (86%), Bengali (82%), Tamil (80%), Aasamease (79%) Punjabi (75%), Odia (72%), and Hindi (71%) form the next tier of top-performing languages.
@@
 These results clearly indicate substantial room for improving recognition performance across Indian languages.
-In the next section we will be evaluating recognition performance in even more challenging end-to-end setting.
-Model Train Data Assamese Bengali English Gujarati Hindi Kannada Malayalam Marathi Odia Punjabi Tamil Telugu Average
-PARSeq [22] Synth 0.47 0.44 0.92 0.32 0.43 0.47 0.52 0.44 0.42 0.4 0.37 0.44 0.47
-BSTD Train 0.79 0.82 0.92 0.61 0.71 0.6 0.58 0.86 0.72 0.75 0.8 0.56 0.73
-CRNN [12] AI4Bharat NA NA 0.27 NA 0.13 NA 0.06 NA NA 0.27 0.16 0.05 0.16
-Off-the-shelf open source models
-Tesseract [37] - 0.15 0.18 0.33 0.12 0.2 0.12 0.03 0.21 0.11 0.19 0.13 0.09 0.15
-PaddleOCR [38] - NA NA 0.61 NA 0.31 0.03 NA 0.36 NA NA 0.28 0.15 0.29
-EasyOCR [39] - 0.12 0.24 0.29 NA 0.18 0.06 NA 0.25 NA NA NA 0.11 0.18
-Table 7: Baseline Results on Cropped Word Scene Text Recognition for 12 languages on BSTD. Please note that PARSeq in used in our proposed baseline – [Uncaptioned image].
-WRR on BSTD-Test
-Method /b Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
-ChatGPT [42] 0.08 0.36 0.09 0.2 0.03 0.05 0.04 0.12 0.21 0.1 0.13
-Google OCR [41] 0.43 0.37 0.32 0.55 0.24 0.37 0.43 0.46 0.52 0.42 0.41
-Tesseract [37] b 0.04 0.02 0.02 0.03 0 0 0.02 0.02 0 0.02 0.02
-SuryaOCR [40] b 0.18 0.15 0.09 0.19 0.07 0.12 0.15 0.21 0.24 0.05 0.14
-[Uncaptioned image] b 0.45 0.37 0.3 0.39 0.29 0.31 0.38 0.41 0.42 0.31 0.36
-[Uncaptioned image] (+oracle TD) b 0.75 0.74 0.54 0.63 0.49 0.54 0.68 0.7 0.69 0.49 0.62
-[Uncaptioned image] (+oracle TD+SI) b 0.8 0.92 0.61 0.74 0.6 0.58 0.72 0.75 0.8 0.56 0.71
-CRR on BSTD-Test
-Method /b Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu Average
-ChatGPT [42] 0.13 0.46 0.15 0.28 0.1 0.14 0.11 0.2 0.31 0.21 0.21
-Google OCR [41] 0.57 0.42 0.36 0.66 0.45 0.55 0.57 0.59 0.69 0.6 0.55
-Tesseract [37] b 0.12 0.07 0.06 0.09 0.05 0.09 0.09 0.09 0.04 0.06 0.08
-SuryaOCR [40] b 0.29 0.24 0.15 0.33 0.19 0.26 0.29 0.31 0.45 0.26 0.28
-[Uncaptioned image] b 0.59 0.47 0.43 0.54 0.48 0.55 0.59 0.53 0.62 0.57 0.54
-[Uncaptioned image] (+oracle TD) b 0.87 0.78 0.64 0.74 0.68 0.8 0.84 0.84 0.78 0.73 0.77
-[Uncaptioned image] (+oracle TD+SI) b 0.93 0.97 0.75 0.89 0.86 0.86 0.89 0.9 0.94 0.83 0.88
-Table 8: We conduct a comparative evaluation of commercial systems, including Google OCR and GPT-4, alongside our proposed baseline namely [Uncaptioned image] for end-to-end scene text recognition on the BSTD dataset.
-The evaluation measures script-wise performance using Word Recognition Rate (WRR) and Character Recognition Rate (CRR).
-Additionally, we present baseline results under two scenarios: (i) when text detection is provided as ground truth (denoted as oracle TD) and (ii) when both text detection and script identification are given as ground truth (denoted as TD+SI).
-“ /b” indicates whether the tool is open-source ( b) or closed-source ().
-Bold and underline number shows best performing results in non-oracle setting for closed and open-source models respectively.
-[Uncaptioned image] is the best open-source model for all the Indian languages.
-The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.
-Method /b Metric Languages Avg.
-Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu
-ChatGPT [42] P 0.11 0.51 0.15 0.28 0.05 0.07 0.06 0.17 0.23 0.11 0.17
-R 0.09 0.42 0.12 0.23 0.05 0.06 0.06 0.15 0.22 0.1 0.15
-F 0.09 0.44 0.13 0.24 0.05 0.06 0.05 0.15 0.22 0.1 0.15
-Google OCR [41] P 0.52 0.54 0.42 0.64 0.34 0.45 0.54 0.59 0.58 0.49 0.51
-R 0.55 0.71 0.44 0.65 0.3 0.42 0.61 0.6 0.64 0.53 0.54
-F 0.52 0.59 0.42 0.63 0.31 0.43 0.56 0.58 0.6 0.5 0.51
-Tesseract [37] b P 0.08 0.05 0.05 0.06 0.01 0 0.04 0.08 0 0.03 0.04
-R 0.08 0.11 0.03 0.05 0.02 0.01 0.04 0.06 0.01 0.02 0.04
-F 0.07 0.06 0.03 0.05 0.01 0.01 0.03 0.05 0 0.02 0.03
-SuryaOCR [40] b P 0.27 0.3 0.19 0.29 0.11 0.17 0.23 0.38 0.34 0.1 0.24
-R 0.24 0.36 0.1 0.23 0.09 0.15 0.18 0.24 0.29 0.13 0.20
-F 0.23 0.29 0.12 0.25 0.09 0.15 0.19 0.28 0.29 0.11 0.2
-[Uncaptioned image] b P 0.59 0.55 0.43 0.53 0.41 0.39 0.47 0.56 0.55 0.38 0.49
-R 0.53 0.46 0.4 0.43 0.33 0.38 0.47 0.49 0.52 0.37 0.44
-F 0.55 0.48 0.4 0.46 0.36 0.38 0.47 0.51 0.52 0.37 0.45
-[Uncaptioned image](+oracle TD) b P 0.72 0.83 0.65 0.66 0.56 0.57 0.65 0.69 0.78 0.54 0.66
-R 0.69 0.75 0.62 0.61 0.5 0.57 0.63 0.68 0.73 0.53 0.63
-F 0.7 0.78 0.62 0.62 0.52 0.57 0.63 0.68 0.75 0.53 0.64
-[Uncaptioned image](+oracle TD+SI) b P 0.75 0.89 0.71 0.71 0.57 0.6 0.66 0.72 0.8 0.55 0.70
-R 0.75 0.9 0.72 0.71 0.56 0.59 0.66 0.73 0.81 0.56 0.70
-F 0.75 0.89 0.71 0.71 0.56 0.6 0.66 0.72 0.8 0.56 0.70
-Table 9: Performance comparison across 10 Indic scripts on the BSTD dataset for end-to-end scene text recognition.
-We report Precision (P), Recall (R), and F1-score (F) for each script and the average performance.
+Futhermore, we show a selection of visual results for several challenging cropped word images from BSTD, comparing predictions from our model against the baseline in Fig. 8.
+Table 10: Comparative evaluation for end-to-end scene text recognition on BSTD using Precision (P), Recall (R) and F-score (F).
+.
 The proposed baseline model, [Uncaptioned image], significantly outperforms available open-source models and achieves competitive performance compared to commercial models.
@@
 The asterisks (*) on Bengali and Devanagari indicate that Assamese and Bengali languages were merged into Bengali, and Hindi and Marathi languages were merged into Devanagari script for evaluation, owing to their shared scripts.
+Method / Metric Languages Avg.
+Bengali* English Gujarati Devanagari* Kannada Malayalam Odia Punjabi Tamil Telugu
+ChatGPT [gpt_4v] P 0.11 0.51 0.15 0.28 0.05 0.07 0.06 0.17 0.23 0.11 0.17
+R 0.09 0.42 0.12 0.23 0.05 0.06 0.06 0.15 0.22 0.1 0.15
+F 0.09 0.44 0.13 0.24 0.05 0.06 0.05 0.15 0.22 0.1 0.15
+Google OCR [googlecloud_visionocr] P 0.52 0.54 0.42 0.64 0.34 0.45 0.54 0.59 0.58 0.49 0.51
+R 0.55 0.71 0.44 0.65 0.3 0.42 0.61 0.6 0.64 0.53 0.54
+F 0.52 0.59 0.42 0.63 0.31 0.43 0.56 0.58 0.6 0.5 0.51
+Tesseract [tesseract] P 0.08 0.05 0.05 0.06 0.01 0 0.04 0.08 0 0.03 0.04
+R 0.08 0.11 0.03 0.05 0.02 0.01 0.04 0.06 0.01 0.02 0.04
+F 0.07 0.06 0.03 0.05 0.01 0.01 0.03 0.05 0 0.02 0.03
+SuryaOCR [paruchuri2025surya] P 0.27 0.3 0.19 0.29 0.11 0.17 0.23 0.38 0.34 0.1 0.24
+R 0.24 0.36 0.1 0.23 0.09 0.15 0.18 0.24 0.29 0.13 0.20
+F 0.23 0.29 0.12 0.25 0.09 0.15 0.19 0.28 0.29 0.11 0.2
+[Uncaptioned image] P 0.59 0.55 0.43 0.53 0.41 0.39 0.47 0.56 0.55 0.38 0.49
+R 0.53 0.46 0.4 0.43 0.33 0.38 0.47 0.49 0.52 0.37 0.44
+F 0.55 0.48 0.4 0.46 0.36 0.38 0.47 0.51 0.52 0.37 0.45
+[Uncaptioned image](+oracle TD) P 0.72 0.83 0.65 0.66 0.56 0.57 0.65 0.69 0.78 0.54 0.66
+R 0.69 0.75 0.62 0.61 0.5 0.57 0.63 0.68 0.73 0.53 0.63
+F 0.7 0.78 0.62 0.62 0.52 0.57 0.63 0.68 0.75 0.53 0.64
+[Uncaptioned image](+oracle TD+SI) P 0.75 0.89 0.71 0.71 0.57 0.6 0.66 0.72 0.8 0.55 0.70
+R 0.75 0.9 0.72 0.71 0.56 0.59 0.66 0.73 0.81 0.56 0.70
+F 0.75 0.89 0.71 0.71 0.56 0.6 0.66 0.72 0.8 0.56 0.70
 4.4 End-to-End Evaluation
-We now present end-to-end evaluation results on the BSTD benchmark. For this, we compare our baseline IndicPhotoOCR with the following methods:
+We now present end-to-end evaluation results on the BSTD benchmark. For this, we compare our baseline with the following methods:
 1.
-Closed-Source Commercial Baselines: We evaluated two commercial systems, namely GPT-4 with vision [42] and Google OCR [41] on the BSTD benchmark [^8] We obtained their results using their best available API on October 31, 2025..
+Closed-Source Commercial Baselines: We evaluated two commercial systems, namely GPT-4 with vision [gpt_4v] and Google OCR [googlecloud_visionocr] on the BSTD benchmark [^9] We obtained their results using their best available API on October 31, 2025..
 Although these comparisons are not entirely fair due to the opaque nature of their training data and model architectures, they offer a useful reference point to gage the current state of the art in Indian language scene text recognition.
@@
 We use the following structured prompt: “Extract all text from this image.
-Return all words in a list format like this: [‘word1’, ‘word2’]”, with a system message defining the model’s role as: “You are an OCR system that extracts text from images.” The interaction was handled through OpenAI’s ChatCompletion.create() API, where both the instruction and the encoded image were passed as part of the message parameter.
-The response of the model was restricted using max tokens = 500.
-Although with more detailed prompt such as also locating text or chain-of-thoughts, this baseline may be improved.
+Return all words in a list format like this: [‘word1’, ‘word2’]”, with a system message defining the model’s role as: “You are an OCR system that extracts text from images.” Although with more detailed prompt, such as asking to locate text or perform chain-of-thoughts reasoning, this baseline may be improved.
 However, such prompt engineering is beyond the scope of this paper.
@@
 (Best viewed in color).
-Figure 11: Failure cases in end-to-end scene text processing: The images showcase representative errors observed across the pipeline.
-(1) Text detection failures from TextBPN++, including false positives, missed detections, merged text, and large text misdetections.
-(2) Script identification errors from the ViT-based model, highlighting misclassified scripts, multi-script confusion, and challenges with degraded text.
-(3) Recognition failures from PARSeq, demonstrating incorrect predictions, missing or hallucinated characters, and cross-script misinterpretations.
-(Best viewed in color).
+Figure 10: A couple of challenging examples with output text by closed source models and ours. Text in red color shows incorrect recognition. (Best viewed in color).
 2.
 Open-Source Baselines: Currently, apart from our own, no open-source system is explicitly designed for Indian language scene text recognition.
-However, we include comparisons with general-purpose OCR tools such as Tesseract [37] and SuryaOCR [40], which are widely used in practical settings.
+However, we include comparisons with general-purpose OCR tools such as Tesseract [tesseract] and SuryaOCR [paruchuri2025surya], which are widely used in practical settings.
 To further analyze the performance of [Uncaptioned image], we also introduce the following oracle variants: (i) IndicPhotoOCR (+Oracle TD): Uses ground-truth text detection.
@@
 These variants help estimate the upper bound on performance in scenarios where text detection or both text detection and script identification are solved reliably.
-Comparison with the aforementioned competitive baselines is presented in Table 8 and Table 9 using WRR/CRR-based metric and Precision-Recall-based metric, respectively.
+Comparison with the aforementioned competitive baselines is presented in Table 9 and Table 10 using WRR,CRR and Precision/Recall and F1 score.
 We observe that [Uncaptioned image] consistently outperforms all existing open-source models across all the evaluated languages.
@@
 Notably, this pattern is also reflected in commercial OCR systems such as Google’s OCR.
+Figure 11: Average Word Recognition Rate (WRR) based on difficulty level. (Best viewed in color).
 We also performed a comprehensive qualitative analysis of the proposed baseline.
 A selection of results is shown in Fig.
-10.
+9.
 We observe that [Uncaptioned image] demonstrates strong multilingual scene text recognition across diverse real-world settings - from street signs and transportation hubs to institutional boards and cultural heritage sites.
-The model performs well even on slanted or perspective-distorted text (e.g., directional signs, railway nameplates), as well as on low-resolution or distant text, indicating robustness in challenging visual conditions.
+The model performs well even on slanted or perspective-distorted text, e.g., directional signs, railway nameplates, as well as on low-resolution or distant text, indicating robustness in challenging visual conditions.
+Further, we show a couple of challenging examples in Fig.
+10 where even closed-source models, including ours, fail to recognize the text correctly.
+Figure 12: Failure cases in end-to-end scene text processing: The images showcase representative errors observed across the pipeline.
+(1) Text detection failures from TextBPN++, including false positives, missed detections, merged text, and large text misdetections.
+(2) Script identification errors from the ViT-based model, highlighting misclassified scripts, multi-script confusion, and challenges with degraded text.
+(3) Recognition failures from PARSeq, demonstrating incorrect predictions, missing or hallucinated characters, and cross-script misinterpretations.
+(Best viewed in color).
 On the BSTD test set, we additionally asked annotators to label each image as easy or hard based on the readability of the scene text.
@@
 Script-wise word recognition rates are presented in Fig.
-12.
+11.
 This clear separation between easy and hard cases highlights the sensitivity of current methods to challenging visual conditions and calls for more robust recognition models.
-Figure 12: Average Word Recognition Rate (WRR) based on difficulty level.(Best viewed in color).
-Figure 13: Recognition errors (words shown in red) from Google Vision, GPT-4, and Refer to caption in comparison with ground truth annotations on images from BSTD test set. (Best viewed in color).
 4.5 Error Analysis
@@
 We performed an extensive error analysis of each module of our pipeline (Section 4.4), summarized in Fig.
-11, and highlight the main failure modes below.
+12, and highlight the main failure modes below.
 Text Detection: Common issues in text detection include (i) false positives where background edges resemble text, (ii) missed detections in low-contrast regions, (iii) merged detections when nearby words are grouped together, and (iv) difficulty with extremely large text, suggesting the model is tuned more for small or medium size words.
-Script Identification: (i) Misclassifications occur among visually similar scripts (e.g.
-Gujarati-Hindi, Bengali-Assamese).
+Script Identification: (i) Misclassifications occur among visually similar scripts, e.g., Gujarati-Devanagari, Bengali-Assamese.
 (ii) Multilingual overlap leads to errors when multiple scripts co-occur, as the model assumes a single script per region.
@@
 These findings emphasize the need for cross-module optimization and feedback-driven refinement between detection, script identification, and recognition.
-In addition, these qualitative analysis, as well as the quantitative analysis in the earlier section, suggest that improving each module in the pipeline can significantly improve overall performance.
-These observations also motivated a visual performance comparison of IndicPhotoOCR with closed-source applications relative to the ground truth, as shown in Fig.
-13.
-5 [Uncaptioned image]: Installation and Usage
-We open-sourced [Uncaptioned image], a toolkit for recognizing text in English and 11 Indian languages: Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Punjabi, Tamil, and Telugu.
-[^9] Future versions will include Urdu and Meitei.
-It is designed to handle the unique scripts and complex layouts of Indian languages.
-By releasing this toolkit publicly, we enable reproducible research, provide a common benchmark for multilingual scene text, and support the development of real-world applications across diverse Indian scripts.
-We provide a user-friendly installation and usage platform, allowing our proposed baseline to be easily replicated.
-We firmly believe that this open-source effort will accelerate future research in Indian language scene text recognition.
-The installation and usage instructions are provided below.
-5.1 Installation
-\$ git clone https://github.com/Bhashini-IITJ/IndicPhotoOCR.git
-\$ cd IndicPhotoOCR
-\$ chmod +x setup.sh
-\$ ./setup.sh
-5.2 Usage
-Detection
->>> from IndicPhotoOCR.ocr import OCR
-# Initialize the OCR system
->>> ocr_system = OCR(verbose=True)
-# Define input image path
->>> image_path = "path/to/image.jpg"
-# Perform text detection
->>> detections = ocr_system.detect(image_path)
-# Save and visualize the detection results
->>> ocr_system.visualize_detection(image_path, detections)
-# Output image saved
-Script Identification
->>> from IndicPhotoOCR.ocr import OCR
-# Initialize the OCR system with automatic language identification model
->>> ocr_system = OCR(identifier_lang="auto")
->>> image_path = "path/to/cropped_image.jpg"
-# Perform script identification
->>> recognized_lang = ocr_system.identify(image_path)
-# Output recognized words
->>> print(recognized_lang)
-Cropped Word Recognition
->>> from IndicPhotoOCR.ocr import OCR
->>> ocr_system = OCR(verbose=True)
->>> image_path = "path/to/cropped_image.jpg"
->>> language = "hindi"
-# Perform text recognition
->>> recognized_text = ocr_system.recognise(image_path, language)
-# Output recognized text
->>> print(recognized_text)
-End-to-End Scene Text Recognition
->>> from IndicPhotoOCR.ocr import OCR
-# Initialize the OCR system with automatic language identification
->>> ocr_system = OCR(identifier_lang="auto")
-# Define input image path
->>> image_path = "path/to/image.jpg"
-# Perform end-to-end OCR
->>> recognized_words = ocr_system.ocr(image_path)
-# Output recognized words
->>> print(recognized_words)
-Detailed documentation for usage is also available here [^10] https://bhashini-iitj.github.io/IndicPhotoOCR/.
+In addition, these qualitative as well as the quantitative analysis in the earlier section, suggest that improving each module in the pipeline can significantly improve overall performance.
 5 Conclusion and Future Work
 In this work, we introduced the Bharat Scene Text Dataset (BSTD), a large-scale benchmark specifically designed to address the gap in Indian language scene text recognition.
-BSTD comprises 6,582 scene images containing 1,06,478 words across 11 Indian languages and English, making it the most comprehensive dataset of its kind.
+BSTD comprises 6,582 scene images containing 1,26,292 words across 11 Indian languages and English, making it the most comprehensive dataset of its kind.
 Through rigorous manual annotation, we ensured high-quality polygon-level bounding boxes and text transcriptions.
@@
 This work is supported by the Ministry of Electronics and Information Technology, Govt. of India, under the NLTM-Bhashini Project.
+Appendix A
+[Uncaptioned image]: Installation and Usage
+We open-sourced [Uncaptioned image], a toolkit for recognizing text in English and 11 Indian languages: Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Punjabi, Tamil, and Telugu.
+[^10] Future versions will include Urdu and Meitei.
+This toolkit is designed to handle the unique scripts and complex layouts of Indian languages.
+By releasing this toolkit publicly, we enable reproducible research, provide a common benchmark for multilingual scene text, and support the development of real-world applications across diverse Indian scripts.
+We provide a user-friendly installation and usage platform, allowing our proposed baseline to be easily replicated.
+We firmly believe that this open-source effort will accelerate future research in Indian language scene text recognition.
+The installation and usage instructions are provided below.
+Installation
+Terminal - Setup
+Usage
+Text Detection Script Identification Cropped Word Recognition End-to-End Scene Text Recognition The tool is reasonably lightweight; the detection module consists of 30M parameters, the script identification module contains 80M parameters, and the 12 recognition modules (23M each) aggregate to 386M parameters in total.
+This architecture offers a viable alternative to large vision-language-based OCR systems, particularly for deployment in resource-constrained and on-device computing environments.
+Detailed documentation for usage of our toolkit is available here [^11] https://bhashini-iitj.github.io/IndicPhotoOCR/.